### v1

# Concept Formation and Alignment in Language Models: Bridging Statistical Patterns in Latent Space to Concept Taxonomy

Abstract
This paper explores the concept formation and alignment within the realm of language models (LMs). We propose a mechanism for identifying concepts and their hierarchical organization within the semantic representations learned by various LMs, encompassing a spectrum from early models like Glove to the transformer-based language models like ALBERT and T5. Our approach leverages the inherent structure present in the semantic embeddings generated by these models to extract a taxonomy of concepts and their hierarchical relationships. This investigation sheds light on how LMs develop conceptual understanding and opens doors to further research to improve their ability to reason and leverage real-world knowledge. We further conducted experiments and observed the possibility of isolating these extracted conceptual representations from the reasoning modules of the transformer-based LMs. The observed concept formation along with the isolation of conceptual representations from the reasoning modules can enable targeted token engineering to open the door for potential applications in knowledge transfer, explainable AI, and the development of more modular and conceptually grounded language models.
1 Introduction
Language models are rapidly approaching human-level language abilities. However, these advancements raise important concerns - on the one hand, despite their immense potential, language models are susceptible to unexpected and potentially harmful behaviors such as hallucination [21], stereotyping [25], misinformation [37], and leakage of sensitive training data [29]. On the other hand, the extent of their capabilities and the nature of their understanding is not yet known, thus capturing the attention of diverse groups, from policymakers to academics [16, 39, 3, 18]. While these issues may seem different, they both stem from a common concern - the "black box" nature. The limited ability to understand how they reach their outputs raises fears about their true capabilities, potential biases, and unintended consequences.
Researchers are actively studying the semantic representations learned by large language models (LLMs), aiming to understand the inner workings of language models. While initially used to study the learning dynamics of linguistic features and abstractions within LMs [50, 26], this research has expanded to explore how these models acquire and represent knowledge about the world. Some studies [38, 18, 1] have explored the ability of transformer-based models to learn representations of color, spatial, and temporal information. In contrast, other studies have raised concerns about the coherence and completeness of this knowledge [16]. The work of [39] even suggests that language models may lack a true conceptual understanding of the real world, echoing the "stochastic parrot" critique [3]. This debate stems from the fundamental question of whether language models merely capture statistical patterns in language or truly grasp the underlying concepts they represent. There are so many works on the semantic relations of the tokens but limited research directly addresses the question of whether these relations are grounded in a conceptual understanding of the real world, or merely reflect co-occurrence patterns in language data [2, 28].
Objectives. To address the aforementioned gap, this paper investigates the process of concept formation in language models. We posit our work based on the argument that ‘Concepts’ are mental representations of categories, which group objects, events, or ideas based on shared features [45]. More generic concepts that correspond to broader categories and share fewer defining features. This is how the hierarchical categories/concepts are formed. In an unsupervised world, when there is no label, categories translate into clusters/communities. Then, similar to [28], we argue that if neural models of language are to have any knowledge about concepts, it ought to be reflected in their token/word embeddings. Thus, communities/categories need to form in the semantic space in which their most relevant category names are often closely related to concepts that describe that category. Note that, while we often form concepts based on categories, not every category will necessarily lead to a clear and distinct concept in the real world. This phenomenon, known as concept alignment [43] will be the focus of our quantitative analysis.
Methodology. Interpretability research traditionally use probing [36, 51]. Probing often involves evaluating models on specific, often artificial tasks designed to isolate particular linguistic phenomena. Probing results might be sensitive to the specific design choices of the probe tasks, such as the choice of dataset, prompts, or metrics, making it difficult to extrapolate conclusions to the model’s overall capabilities in real-world scenarios. Thus, while effective for specific domains, they have inherently limited ability to provide broader insights into the extent of models’ underlying representations and reasoning capabilities beyond specific tasks they are tuned for [48] Therefore, we avoid using conventional probing, instead, we propose a concept extraction methodology based on graph creation (K-nearest neighbors), fuzzy weighting mechanism (borrowed from UMAP category theory [32]) and community/cluster detection algorithms to identify groups of highly interconnected data points, representing potential concepts. The performance of this approach hinges upon the existence of a meaningful, hierarchical structure within the input embedding layer. This method can potentially offer a deeper and more holistic understanding of the underlying knowledge encoded in these representations (as compared to probing).
Models Under Investigation. Using our proposed method, we investigate the embedding spaces of GloVe, a traditional statistical language model [40]; ALBERT, a transformer-based encoder language model [23]; and T5, a transformer-based autoregressive language model [42] to comprehend if concept formation occurs in any of these models.
Evaluation. In order to answer the question of "if concept formation happens in the semantic space," we evaluated the formed clusters against external references such as WordNet [34], Name Dataset [44], and country-state-city databases [11] for most of our evaluations. For clusters where open-source datasets were not available, we employed human annotators as an external reference. We observed that the process of concept formation can occur as early as the input embedding layer in transformer-based models. Considering that this observation hints at the existence of semantic memory in transformer-based models, we further investigate in Section 4.3 whether there is a separation between the identified semantic memory and the reasoning mechanisms. Specifically, we ask whether the model can still reason effectively with a modified semantic memory. This line of inquiry aims to deepen our understanding of the relationship between knowledge representation and reasoning capabilities within transformer-based models.
Figure 1: Simplified steps on how external information is understood and retained. Upon understanding a newly encountered word/entity, it is typically stored in the semantic memory. The existence of semantic memory (on the left) allows the previously encountered words/entities to have a form of meaning even without requiring an external context. The scatter box on the right is the community (primarily associated with moving creatures) we extracted from the ALBERT model [23].
2 Semantic Representation Space
2.1 Concept formation, Representation Learning and Semantic Memory
A simplified model of human understanding can be described as a multi-step process in which the incoming sensory information is integrated and turned into a brain-constructed interpretation of external objects/stimuli. This results in a mental construct (or percept) [15]. Then, the newly formed mental construct gets compared, integrated, and associated with the existing knowledge already retained in the semantic memory. These connections help us to conceptualize and understand the new information based on what we already know (as shown in Figure 1). While diverse reasoning mechanisms exist, a common thread among them seems to be their reliance on and interaction with the semantic memory. This interaction likely leaves enduring traces within such a memory [5]. The integration of newly encountered mental constructs within an existing cognitive framework is often guided by their relations to established internal constructs [13]. This process facilitates the creation, refinement, or expansion of semantically similar clusters [27], implying a degree of inherent categorization [^1] This notion implies an interrelation between the recognition and categorization..
Representation Learning and Concept Formation Representation learning is a fundamental aspect of language models, where the goal is to learn distributed representations (embeddings) for words or subword units that capture their semantic and syntactic relationships [4]. Early models like word2vec [33] utilized shallow neural networks to learn these embeddings from large text corpora. More recent models, such as Large Language Models (LLMs) like GPT-3 [30], employ transformer architectures with self-attention mechanisms [53], enabling more accurate and dynamic representation learning. From the distributional hypothesis, models can form "concepts" by identifying patterns and relationships within data, particularly through recognizing the approximate invariance of shared features across different data instances [13]. It should be noted that "concept formation" is a byproduct of the learning process, not a guaranteed outcome. Furthermore, the formed "concepts" might not always align with human-defined concepts. Thus, it is crucial to investigate and evaluate the nature of concept formation within models and their alignment with human understanding.
2.2 Evaluation Methods
The analysis, measurement, and interpretability of semantic representation learning in language models have been the subject of extensive research. Various methods have been proposed to evaluate how well these models capture semantic meaning.
Intrinsic evaluation methods, assess semantic representations by measuring word embedding similarity, comparing them to human judgments of relatedness, (e.g. [33, 17, 52, 36, 51]). While these methods provide valuable insights, they often suffer from the limitations that probing techniques impose, and may not fully capture how well a model forms conceptual hierarchies or grasps the full spectrum of semantic relationships, as word embeddings typically represent individual words rather than higher-level categories [9]. This limitation can hinder the evaluation of a model’s ability to understand broader semantic relationships and its capacity for abstract reasoning.
Another approach is to use extrinsic evaluation methods, which measure the performance of language models on downstream tasks that rely on semantic understanding, such as sentiment analysis, question answering, and machine translation [54, 55]. The performance on these tasks can indirectly indicate the quality of the learned semantic representations. The main challenge with using such extrinsic evaluation metrics to assess concept formation in the embedding space is that they do not directly measure the quality of the concepts themselves. Instead, they measure how well the model performs on specific tasks that rely on those concepts, which does not necessarily prove that the model has formed robust, human-like concepts. For instance, a model might accurately classify sentiment without truly grasping the nuances of emotions like sarcasm or irony [31].
In addition to evaluation metrics, researchers have also employed visualization methods to interpret learned representations. These methods often involve projecting high-dimensional embeddings into lower-dimensional spaces for visualization [48, 46]. However, several challenges arise with these approaches. Firstly, they offer indirect and subjective assessments of model understanding, lacking a quantitative basis for evaluation. Secondly, the dimensionality reduction process can lead to information loss [14, 56], potentially obscuring important nuances in the representations. Finally, different visualization techniques can produce conflicting results, making it difficult to reach definitive conclusions about the model’s true comprehension. To mitigate these shortcomings, we propose a concept extraction mechanism that identifies communities in higher-dimensional space. This approach enables both quantitative evaluation and mitigation of potential information loss incurred during dimensionality reduction.
3 Concept Extraction
As discussed in Section 2.1, identifying clusters or communities within the semantic representation space is a crucial initial step in evaluating concept formation. We prioritize community detection among clustering methods for the following reasons:
  • Alignment with Semantic Network Structure: The underlying semantic representation, as detailed in Section 2.1, exhibits a network-like structure. Community detection algorithms are specifically designed to uncover such inherent community structures within networks, making them better suited to our task.
  • Unknown Number of Clusters: Traditional clustering algorithms are often sensitive to the choice of the number of clusters or other clustering parameters [57]. In our case, we lack this prior knowledge and only assume that semantically similar words are closer in the representation space. Community detection methods do not necessitate a predetermined number of clusters.
  • High Dimensionality of Data: The high dimensionality of the semantic representation space poses challenges for many clustering algorithms [57]. Community detection methods on the other hand, naturally circumvent this issue by transforming the high-dimensional space into a graph , where relationships between data points become more readily apparent.
This approach allows for a more effective and accurate identification of semantic clusters, reflecting the natural structure of the semantic representation space and facilitating a deeper understanding of concept formation.
3.1 Graph Construction
While categories themselves are not strict clusters, they may exhibit internal structure or fuzzy partitions based on shared features or relationships. Considering our discussion in the previous section, we seek a representation space where the structure of these categories is explicitly reflected in the geometry of the manifold. This requirement arises from the desire to directly manipulate and modify these clusters through geometric operations on the representation, opening up possibilities for Token Engineering within the language models.
Inspired by the effectiveness of the UMAP dimensionality reduction [32], we use the Nearest neighbor descent to find the $K$ nearest neighbors for every token embedding in the embedding space and then use the following equations as the weight function of the edge between $x_{i}$ and $x_{j}$ nodes in the $K$-NN graph [32]:
$\omega(x_{i},x_{j})=exp\left(\frac{-\max(0,~{}d(x_{i},x_{j})-\rho_{i})}{\sigma_{i}}\right)$ (1)
where $d(\cdot,\cdot)$ is the distance function (cosine in our case) and $\rho_{i}$ is calculated as:
$\rho_{i}=\min\{d(x_{i},x_{j})\mid 1\leq j\leq k,\;d(x_{i},x_{j})>0\}$ (2)
where $k$ is the number of neighbors of node $i$. Finally, $\sigma_{i}$ is calculated by setting the summation of weights of a node to be equal to a constant (i.e., $\log_{2}(k)$):
$\sum_{j=1}^{k}\omega(x_{i},x_{j})=\log_{2}(k).$ (3)
Building upon our theoretical arguments, the invariant representations (if they exist), should form fuzzy partitions that can be detected by the community detection algorithms. Thus, for the third step, we perform community detection.
3.2 Louvain Community Detection
The Louvain method is a widely used algorithm for community detection in large networks. It finds the communities by optimizing a metric called modularity. The modularity of a partition is a scalar value between -1 and 1 that measures how much more densely connected the nodes within a community are compared to how connected they would be in a random network. [6]. For a weighted graph, modularity is defined as:
$Q=\frac{1}{2m}\sum_{i,j}\left[A_{ij}-\frac{k_{i}k_{j}}{2m}\right]\delta(c_{i},c_{j})$ (4)
where $A_{ij}$ represents the weight of the edge between $i$ and j; $k_{i}$ is the sum of the weights of the edges attached to vertex $i$; $m$ is the sum of all of the edge weights in the graph; the $\delta$-function $\delta(c_{i},c_{i})$ is 1 if $u=v$ and 0 otherwise; $c_{i}$ is the community to which the nodes $i$ belongs to.
Then, it aggregates the communities to identify hierarchical structures. In this phase, each community is considered as a single node and the links between the new nodes are calculated as the sum of the weight of the links between nodes in the corresponding two communities. More details are given in Algorithm 1.
Data: The initial input is a weighted network of all the nodes exist in the entire space.
Result: A set of hierarchical communities.
1 Community Detection;
2 Create a community list; assign a different community id to each node of the network;
3 while a local maxima of the modularity is not attained do
4       for each node i do
5             for each neighbor j do
6                   evaluate the gain of modularity if i moved to the community of node j;
7                   keep the maximum gain and community id;
8
9            if the maximum gain is positive then
10                  Move node i to the community with maximum gain.
11
12
13Community Aggregation;
14 if Number of Communities > 1 then
15       Reduce each community to a single node;
16       go to 1
Algorithm 1 Louvain
3.3 Concept Extraction Algorithm
for our concept extraction algorithm, as the first step, we create and weight the adjacency graph using $K$-NN, UMAP-based weighing formula (mentioned in section 3.1), and then use Louvain [6] and label propagation [58] community detection algorithms.
There are multiple viable strategies to extract hierarchical communities in our methodology; the first one is to use algorithms such as Louvain that inherently generate hierarchical communities [6]. However, the summation of the weights as a new weight in the community aggregation phase of Louvain algorithm skews the weighted graph in favor of merging smaller communities in the next phase. This detaches the community detection from the actual values in the semantic representation space (i.e. graph weights higher in the Louvain hierarchy no longer reflect the geometrical affinity of the nodes). Thus, for our concept extraction algorithm, we use 1-2 community aggregation and rather use KNN iteratively with different granularity for extracting hierarchical concepts (more details are described in Algorithm 2).
Note that we only use well-established methods such as $K$-NN, UMAP-based weighing formula, as well as Louvain [6] and label propagation [58] community detection algorithms to capitalize on the established generalizability of these algorithms. Although, as we show in the next sections that our method produces amazingly good categories, it should be noted that we intend to focus our analysis on “if the language model forms concepts" rather than creating the most optimal concept extraction mechanism. Thus, as an extension to this work, one can focus on further optimizing our proposed method. Notably, our concept extraction is algorithm-agnostic; alternatives could be readily employed.
Data: All tokens in the input embeddings.
Result: A set of hierarchical communities.
1 Create a community list. The initial community is the entire space;
2 for k=[different neighbor sizes] do
3       for all communities do
4             (1) Generate KNN graph of the community using the input embedding weights;
5             (2) Compute the edge weights of the graph using fuzzy simplex;
6             (3) Apply Louvain/Label-Propagation community detection;
7             (4) Add the identified communities to the list;
8
9
Algorithm 2 Concept Extraction
3.4 Models Under Investigation
Using our proposed method, we investigated the embedding spaces of GloVe, ALBERT, and T5 to determine if concept formation occurs consistently across these diverse language models. We chose this spectrum of models to: (i) assess if our method for identifying concept formation is applicable across a wide range of architectures and training data, and (2) investigate whether the concept formation is a fundamental property of language representation, regardless of the specific model architecture or training data. We hypothesize that the language model with the least embedding layer dimension is more appropriate for our investigative analysis, because, when all other factors are held constant, models with more compact latent spaces generally produce more invariant representations [20]. Thus, for the detailed analysis and evaluations, we pick ALBERT  [^2] We used open source Hugging Face albert-base-v2 and albert-xxlarge-v2 (https://huggingface.co/albert). [23] as our primary target since it has the tightest latent space dimensionality.
Finally, we hypothesize that exploring the formation of such analogs early in the language model, potentially at the input embedding layer, could open the door to the possibility of knowledge-reasoning separation. Thus, we focus our investigation on this layer. More details about the models can be found in Appendix A.
4 Experiments and Discussions
Concept formation involves the creation of abstract internal representations through the clustering of inputs with shared features. While numerous intrinsic clustering metrics exist to assess cluster formation quality, their application to our use case is limited by two factors:
  • The formed clusters are not situated within Euclidean space [24, 8, 7], rendering geometric properties such as cluster distances inadequate indicators of concept well-formedness and distinctiveness.
  • Concepts inherently possess a degree of vagueness [19], thus metrics like compactness or separation do not reliably reflect the quality of formed concepts.
Consequently, to assess the quality of the formed concepts, we shift our focus to evaluating concept alignment, specifically examining both LM-LM and Human-LM alignment. LM-LM alignment allows for the generalization of our observations to other language models, while Human-LM alignment serves as an indicator of the quality and real-world interpretability of the formed concepts.
Concept Extraction Setup. We configured our algorithms to create $k$-NN graph iteratively for different values of $k$. This enables us to observe the communities/concepts formation at various granularities. Table 1 shows the number of identified clusters for $k=[100,75,50,25,12,6]$.
Table 1: Number of communities with different granularities of nearest neighbors for ALBER, T5 and Glove. For Glove, we only used the subset of GloVe that present in Albert Vocab
$K$ settings Vocab
Models 100 75 50 25 12 6 Size
T5 1 115 1203 4551 8137 9407 32000
Albert 8 133 1058 4442 7718 8626 30000
GloVe 9 207 1157 3521 6237 7200 25869
4.1 LM-LM alignment
To evaluate LM-LM alignment, we measure the community mapping across the language models under test. We define a mapping as the ratio of the successfully mapped communities from the embedding space of the language model with the smaller vocabulary to the language model with the larger vocabulary. Furthermore, we define success in mapping process as if the ratio of of found tokens from source community in the destination community is greater than a certain threshold. thresholds greater than 0.5 imply at worst many-to-one mapping (when the number of tokens in the source and destination space are similar, it is more likely to imply one-to-one mapping). Such one-to-one mapping implies a direct concept alignment, and many-to-one mapping implies the source community is aligned with hierarchies of the destination concept.
This mapping score provides a quantitative measure of the degree to which the concept communities formed by one model can be translated to the other. To refine our analysis, we filtered the tokens to include only those that start with whitespace in T5 and ALBERT models. Empirically, this enhances the likelihood of associating the token with a word existing in the external world, thereby enhancing the validity of our concept alignment assessment. Note that to find the mappings we only used the token sets that are shared between the model pairs. That is 25869 for Albert/GloVe, 14861 for T5/GloVe and 5055 tokens for T5/Albert pairs.
Table 2: Pair-wise Mapping score of the formed communities in the Glove, Albert, and T5 embedding spaces.
Thre- Albert / GloVe T5 / GloVe T5 / Albert
shold $K$=75 $K$=50 $K$=25 $K$=12 $K$=6 $K$=75 $K$=50 $K$=25 $K$=12 $K$=6 $K$=75 $K$=50 $K$=25 $K$=12 $K$=6
0.5 0.70 0.57 0.62 0.68 0.70 0.62 0.74 0.84 0.89 0.91 0.61 0.78 0.91 0.95 0.97
0.6 0.56 0.42 0.43 0.42 0.42 0.46 0.60 0.70 0.72 0.72 0.43 0.60 0.80 0.86 0.86
0.7 0.42 0.30 0.30 0.28 0.27 0.27 0.46 0.54 0.56 0.56 0.30 0.49 0.70 0.79 0.80
0.8 0.32 0.22 0.24 0.23 0.21 0.15 0.39 0.48 0.52 0.52 0.24 0.45 0.69 0.78 0.79
0.9 0.17 0.16 0.20 0.21 0.21 0.07 0.32 0.45 0.50 0.51 0.18 0.43 0.67 0.78 0.79
[…]
The aforementioned observations strongly support the generalization of LM-LM concept alignment. The high degree of alignment, particularly for granular concepts and between models with similar architectures (like ALBERT and T5), indicates a shared understanding of semantic relationships across these models. This opens up exciting possibilities for leveraging this alignment to improve model performance or transfer knowledge between models without the need for extensive retraining. However, further research is needed to explore the extent and limitations of this potential knowledge transfer, especially for more abstract concepts and between models with different architectures.
Case Sensitivity Analysis. The T5 model’s vocabulary preserves case information, enabling us to examine how formed concepts align with capitalization differences. We identified 4,328 tokens with varying case appearances (total of 8,887 tokens). We found that for highly granular concepts (k=6), 80% of these tokens belong to the same community. This ratio increases to 85% for k=25 before plateauing, suggesting that case variations generally do not significantly alter the semantic grouping of tokens. This finding supports the notion that the model learns to associate words with their meanings regardless of capitalization, particularly for more abstract or broader concepts (larger k values). However, the initial increase in alignment ratio with increasing k implies that case sensitivity might still play a minor role in differentiating highly specific or nuanced concepts.
4.2 Human-LM alignment
In the previous section, we established the concepts formed in LM models have a degree of alignment with respect to each other. To investigate human-LM alignment, in this section, we focus our analysis on the ALBERT model. While our empirical results are derived from ALBERT, the principles discussed are generalizable to other LLMs, given the LM-LM alignment established earlier (we include human-LM alignment evaluations for GloVe and T5 in Appendix B).
Concept Hierarchy Formation. When examining the broader perspective (i.e., $k$=100), the model primarily found groups of named entities (mostly names of people and places), adverbs, sub-words, some number symbols, and a super-community for ‘other vocabulary’ (appendix C, Figure 4 shows the overall concept hierarchies of the ALBERT vocab). Zooming in further (e.g., $k$=75), these communities revealed more specific clusters that are relatable to the real world. For instance, within named entities, clusters formed for personal vs. location names, even further pinpointing locations by country. Similarly, within the aforementioned super-community, sub-clusters emerged that mapped to various lifestyle aspects, including relationships, social structures, multimedia, and motile creatures. As the granularity level increases (approaching a smaller $k$ value), clusters exhibit a stronger association with word forms.
Concept Alignments. To investigate the presence and extent of human-LM concept alignment, we systematically examine the model’s representations of distinct facets of the external world:
  • Numerical Concepts. This category focuses on numerical values, facilitating a targeted analysis of intra-category relationships, such as value-based ordering.
  • Named Entities (Names and Locations). Primarily comprising proper nouns, this category leverages readily available labeled datasets to validate alignment between the LM’s conceptual groupings and established knowledge bases. It also allows us to investigate if associative relations (e.g., country-nationality) influence community formation.
  • Social Structures. This category reflects the LM’s approach to forming categories that encompass both abstract and concrete nouns, including roles, titles, and processes. This analysis can elucidate its ability to grasp and align complex relationships and concepts.
By analyzing these diverse facets, we aim to comprehensively evaluate the degree to which the LM’s internal representations align with human understanding of the external world.
Table 3: Orderliness score of different communities of the numbers. Years refer to the cluster of numbers between 1816 - 2021. Support refers to the number of samples within the cluster.
Category Top-1 Top-3 Top-5 Support
0-100 0.68 0.89 0.92 100
100-262 0.75 0.94 1.0 115
263-300 0.46 0.78 0.89 37
300-400 0.42 0.71 0.84 82
400+ 0.53 0.75 0.83 110
Years 0.86 0.96 0.98 203
4.2.1 Symbols-Numbers
At a high level, numbers and symbols are completely separated from the rest of the tokens in the embedding space. within themselves they form communities according to (i) years, (ii) integer values, (iii) tokens indicating monetary values (e.g., \$1), (iv) ratio/time (e.g., 2:1, 3:30), (v) fractions (e.g., ’/8’), (vi) large values (e.g., ‘100,000’), and (vii) percentages (e.g., $42\%$). Within these communities, integers form communities based on their hundreds. Within the embedding manifold, we analyzed the spatial distribution of number sequences, particularly focusing on their sequencing. Since the embedding manifold is not guaranteed to be a Euclidean structure [24, 8, 7], conventional distance-based measures of order were not applicable. We define a point $x$ as "locally ordered" on a manifold if it resides within the intersection of the $k$-nearest neighbor sets of both its immediate neighbors, $x+1$ and $x-1$. This localized approach captures the intuitive notion of orderliness within a small neighborhood on the manifold, rather than relying on a global ranking.
Orderliness Score. Table 3 presents the orderliness scores of the communities identified by our algorithm. The "top-1" score represents strict orderliness, where $x+1$ lies exactly between $x$ and $x+2$ in the embedding space. The "top-3" and "top-5" scores measure relaxed orderliness, where $x+1$ falls within the 3 or 5 nearest neighbors of both $x$ and $x+2$, respectively. As the results indicate, all detected communities exhibit a higher degree of local orderliness, regardless of the chosen level of strictness. This finding is significant because it suggests that language models may possess not only the ability to categorize heterogeneous input entities but also the capacity to construct meaning within smaller, internally consistent structures (internally homogeneous sub-structures). This implies the potential for LMs to move beyond simply classifying information to actively interpreting and generating meaning within specific contexts.
Figure 2: Visualization of the identified name and location communities of size larger than 10 entities. UMAP projection as well as kernel density estimate (KDE) plot is used for the visualization.
4.2.2 Names and Locations
Figure 2 visualizes the major identified communities of locations and human names. Generally, the formed clusters associate with specific regions or cultures and contain both location and personal names. Within these, even more granular sub-clusters emerge, characterized by distinct communities of location and personal names. Interestingly, we observed a degree of geographic ordering within the identified communities. As illustrated in Figure 2, there appears to be a general trend from east to west as we move across the communities from bottom-left to top-right. The leftmost communities are predominantly associated with Japanese locations and names, while those on the rightmost side are primarily linked to Europe and the United States. This suggests that the model’s internal representations in the input embedding layer may inherently capture geographical relationships.
To mitigate the subjectivity risk of assessing the semantic structure, we further used external datasets in our evaluations. For Names, we used name-dataset [44] which consists of a comprehensive set of names (730K first names and 983K last names), their associated genders, and their popularity rank for each country. For locations, we used the country-state-city database [11] which contains information on 250 countries, 5K+ states, and 150K+ cities. Table 4(a) shows the high-level communities that our approach detected. Most of the high-level communities are a mix of names/locations associated with certain geographical/cultural regions. Within these clusters, names and locations form distinct subclusters which we discuss in more detail in the following subsections.
Table 4: Evaluation of the location and name tokens for different hierarchy levels. 4(a) presents the precision for high-level communities, while 4(b) and 4(c) show the more granular communities formed within the detected communities in 4(a)
(a) Precision of the largest identified name and location communities with respect to name and location databases. Note that these are at a higher level in the cluster hierarchies. Table 4(b) shows the identified granular sub-clusters and their associated precision.
Category Precision Support Note
US/UK/AUS/NZ 0.882 1011 Human & Location Names
Male 0.854 946 Human Names
Female 0.866 552 Human Names
West-Asia 0.684 390 Human & Location Names
Hispanic/Latino 0.685 282 Human & Location Names
US 0.720 267 Location Names
Europe 0.739 215 Human & Location Names
East-Asia 0.741 178 Human & Location Names
(b) The precision of communities based on the identified categories with respect to name-database. First indicate categories of first names last indicates last names
Country Overall Country Sup- Note
Precision Precision port
USA/UK 0.857 0.725 211 First
UK/Canada 0.886 0.698 116 First
Saudi/Arabic 0.82 0.76 94 First
Spain/Mexico 0.977 0.78 87 First
USA/Hebrew 0.835 0.568 81 First
Italy/Swiss 0.887 0.625 80 First
Belgium/France 0.928 0.789 76 First
German/Sweden 1.0 0.709 55 First
German/Austria 0.962 0.717 53 First
India 0.82 0.56 39 First
Russian 0.896 0.724 29 First
France 0.9 0.737 76 Last
Mexico 0.951 0.855 83 Last
China 0.9 0.9 30 Last
Denmark 0.88 0.64 25 Last
Germany 0.95 0.95 20 Last
Japan 0.928 0.857 14 Last
(c) The precision of communities based on the identified categories with respect to the location database.
Country-Region Precision Support
United States 0.80 240
Germany 0.4125 80
France 0.409 66
Africa 0.69 55
India 0.58 50
Italy 0.590 44
Mexico 0.424 33
Spain 0.592 27
China 0.5 20
Japan 0.736 19
Philippines 0.46 15
Pakistan 0.461 13
Netherlands 0.636 11
North-Africa 0.8 10
Names. To determine whether the sub-clusters are associated with names, we pre-filtered the clusters that at least 70% of their tokens are in the top 1000 names (of any country), with gender confidence of above 0.8. We primarily use gender confidence to distinguish between first-name from last name clusters. Then we cross-referenced all identified name communities for all countries in the dataset and assigned the country name with the highest score  [^3] For some clusters we picked the second highest country name if the scores were similar. Due to the cosmopolitan nature of countries like the USA, They tend result in a high score across the board. as the cluster names.
Table 4(b) shows the precision score of the identified granular communities. The first column shows the ratio of the community members that are indeed human names (overall precision), while the second column shows the ratio with respect to specific countries. It should be noted that the country-wise scores of the countries with similar cultures and languages were similar. The reported precision score shows a high degree of categorization based on the country/culture of origin.
It should be noted that we only included clusters with sizes larger than 10 and country-wise precision of larger than 0.5 due to space limitation. The list of identified cluster names goes far beyond the aforementioned table. Clusters such as character names from books, mythology, and car brand names were also identified which were not included due to space limitations.
Locations. It should be noted that, although the dataset is somewhat comprehensive, the LM token space associates only one token to each location, leading to an artificial decrease in precision for multi-token location names. Despite these limitations, we were able to identify communities within the input embedding space that are associated not only with the location category but also with specific regions/countries. The precision numbers in Table 4(c) suggest the model groups the locations with respect to the borders of the countries, which, to a certain degree, implies a subjective perception of geographical knowledge aligned with human perception (since it is evaluated against human annotated dataset). This finding suggests that the language model attempts to construct a human-like mental map  [^4] We refer the definition of the mental map to the [American Psychology Association dictionary] (https://dictionary.apa.org/mental-map). It is defined as “a mental representation of the world or some part of it based on subjective perceptions rather than objective geographical knowledge., wherein it approximates borders and associates nationalities with specific clusters/communities. It should be noted that, due to space constraints, we only present the communities containing more than 10 entities  [^5] The complete set of clusters is included as the supplementary material.. Note that dealing with multi-token location names is more challenging. For instance, "Carolina" was correctly clustered within the United States community by the LM, while our reference dataset misclassified it as a city in Brazil. Additionally, not all location names with English spellings are included in the dataset; for example, ’Wurttemberg’ (or ’Nuremberg’), a region within Germany, is absent, leading to an artificial decrease in precision.
4.2.3 Social Structures
Our methodology identifies a cluster of 903 members with the theme of social structures. In order to have a reference dataset, we annotated the dataset using a combination of GPT-4 and human annotators 666 A human annotator was used alongside GPT-4 to correct GPT-4 misclassifications. Label correction by a human happens faster than the label suggestion task. We estimated our approach is more time/cost-effective while resulting in the same quality annotations.. Since analyzing such concepts can accompany a degree of human subjectivity, we feed the datasets to GPT-4 and ask to identify the theme of each category (as shown in Table 5) for mitigating such uncertainity. In order to reduce subjectivity (we formed the class names based on the GPT-4 recommendation. Then, we asked GPT-4 and human annotator to classify each word with respect to given class names (we added another class named “Other" to avoid forcing the annotators to misclassification).
Then, we calculated the precision score of the identified clusters with respect to our annotated dataset. We see the precision scores shown in Table 5 as evidence that the model forms an idea on different aspects of social structure in its semantic memory. When it comes to more granular clusters ($k$=25), the sub-clusters are mostly word-forms or semantically similar words.
Intriguingly, the vocabulary model grouped words like "God" and "divinity" within the same community as concepts and structures associated with Christianity. Conversely, terms like "Islam," "Judaism," and "Talmud" formed a distinct cluster. This finding warrants further investigation to determine whether it reflects potential biases within the underlying semantic memory.
Table 5: The precision score of community members belonging to the identified categories.
Category Precision Support
Religious (Christianity) 0.818 258
Military and Law Enforcement 0.842 133
Administration and business 0.788 129
Political Ideologies/Movements 0.648 125
Monarchy and Aristocracy 0.64 107
Legislature and Election 0.736 80
Figure 3: Visualization of the social structure cluster and its associated identified sub-clusters.
4.3 Knowledge-Reasoning Separation
Thus far, we have demonstrated that the LM constructs a knowledge base (mental representation) directly within its input embedding layer. Furthermore, we have established a degree of human-LM alignment in both hierarchical structure and semantic meaning.
Table 6: Huggingface ALBERT base model on GLUE, SuperGLUE, and SQUAD tasks. For the baseline, the model was finetuned without altering the embeddings. For the mid-point, the embedding layer entries are assigned the mid-point embedding of their associated community. We used [41] repo for benchmarking.
mid-point GLUE
Method/Tasks mnli mrpc qnli qqp rte sst stsb wnli
baseline 0.827 0.841 0.902 0.858 0.765 0.915 0.872 0.549
mid-point 0.849 0.865 0.910 0.875 0.783 0.922 0.890 0.563
mid-point SuperGLUE SQUAD SQUAD
Method/Tasks boolq cb copa multirc record wic wsc v1 (f1) v2 (f1)
baseline 0.622 0.512 0.59 0.350 0.586 0.595 0.528 83.72 70.9
mid-point 0.621 0.478 0.55 0.372 0.588 0.626 0.634 84.4 74.9
Now, we are interested to see the extent to which the knowledge learned during pretraining is modular and separable from the reasoning mechanisms employed by the language model in downstream tasks. For example, can the knowledge learned during the pretraining phase be selectively removed or modified without significantly impacting the model’s performance on finetuning? The modularity can also impact the effectiveness of Language Model Inversion [35, 49] techniques, which aim to extract private information such as names or other sensitive details learned during the pre-training777 For example, if knowledge is found to be highly modular, it may be possible to develop targeted interventions that obscure or remove specific sensitive information without significantly impacting the model’s overall performance on downstream tasks..
To investigate this, we selected GLUE, SuperGLUE, and SQUAD benchmarks as downstream tasks to assess language model performance. We then systematically removed within-community information by calculating and assigning the embedding space center (mid-point value) of each community to all its members. For example, in a community of names like "James," "John," and "Alex," all members would share the same embedding. Table 6 shows the results on major LM benchmarks GLUE, SuperGLUE, and SQUAD for the ALBERT base model. Although our experiment does not prove the separation of knowledge and reasoning, it indicates that at least the granular information acquired during the pretraining is not required for the model’s performance on the aforementioned LM benchmarks. This is significant because it opens the door for embedding engineering of private or harmful information that is learned during the pretraining.
5 Limitations and Risks
Limitations: The model forms the conceptual communities that are meaningful but its priority is not exactly the same as that of humans. The model vocabulary is a contributing factor to the way the model prioritizes the formation of conceptual clusters in its embedding layer. For example, the number of English names is much higher than the other languages and this has caused the model to form high-level communities (e.g., $k$=75) specified for names vs. less frequent names/locations a high-level community contains the combination of regions personal and location names. This limits our method to associate the KNN resolution with the abstraction level of the extracted concepts/categories.
Risks: This work provides detailed information about (1) the formed clusters/concept in the input embedding layer, and (2) the separation of memory from reasoning in ALBERT. As the methodology can also be applied to other models, it can potentially facilitate more advanced adversarial attacks and content manipulation in language models.
6 Conclusion and Future Works
This research focused on concept formation and alignment across different language models (LMs) and between LMs and humans. We propose a concept extraction mechanism that uncovers the emergence of distinct conceptual communities within the entire embedding space. This approach gives a more holistic view of conceptual representation than examining specific samples, as it explores the full landscape of concepts embedded within the models.
We examined GloVe, ALBERT, and T5 LMs and found that all of them form communities that can be aligned with the external world categories/concepts. We further investigated various aspects of the emergent conceptual communities, ranging from their numerical sequencing to the interrelationships and associations found within the blend of concrete and abstract concepts. Across various language models and types of concepts, distinct communities emerge in the embedding space. These communities demonstrate a significant alignment with human-understandable concepts, suggesting that statistical patterns, rather than specific architectures, or conceptual domains, drive concept formation across LMs. However, the quality and alignment of the formed concepts is a function of the architectural choices, training data, and tokenization schemas.
We further observed that concept formation in transformer-based LMs occurs as early as the input embedding space. This suggests that the input embeddings themselves encode substantial semantic information, contrary to traditional views where meaning emerges later. Moreover, the emergent concepts within the input embedding layer exhibit a degree of separation from subsequent transformer layers.
Early concept formation has several implications, opening exciting avenues for future line of works:
  • Knowledge transfer across tasks and models: By identifying and leveraging shared conceptual structures across different tasks and models, early concept formation can enable more efficient transfer learning and reduce the need for extensive retraining on new tasks.
  • Improving model architectures, training, and evaluation: Understanding how concepts emerge early in the model can inform the design of architectures that better capture and represent meaning, leading to improved performance and interpretability. This knowledge can also guide the development of more effective training techniques and evaluation metrics.
  • Mitigating training data information leakages and reducing biases: Early concept formation can be utilized to identify potential biases and information leakages in the input embedding layer. This opens the door to the development of token embedding engineering techniques to mitigate these biases without requiring re-train.
Appendix A Models Under Investigation
A.1 ALBERT
ALBERT (A Lite BERT) [23] is a transformer-based model for language representation learning, designed to be more efficient than its predecessor, BERT [10]. While it shares the same basic architecture as BERT, it incorporates two main key modifications:
  • Factorized embedding parameterization. The benefit of factorized embedding parameterization in ALBERT is the significant reduction in the number of parameters compared to models like BERT. In BERT, the word embedding size (E) is tied to the hidden layer size (H), leading to a large embedding matrix as H increases. ALBERT instead factorizes this embedding into two smaller matrices, one projecting token ID vectors to a lower-dimensional space (E) and another projecting from this space to the hidden layer (H). This allows H to be much larger than E without increasing the parameter count of the embedding layer substantially, resulting in a more efficient use of parameters. This is particularly beneficial for large models, where memory limitations can hinder training and deployment.
  • Cross-layer parameter sharing. Parameter sharing acts as a form of regularization, preventing the model from overfitting to specific layers or features in the data. This can lead to improved generalization performance on unseen data.
Note that reducing the number of parameters and sharing information across layers can force the model to learn more general representations, thus indirectly contributing to better concept formation. datasets. ALBERT is pretrained on English text datasets, namely the English Wikipedia and BookCorpus, using self-supervised learning objectives.
Tokenization. It uses Sentencepiece tokenizer [22] on the uncased corpus with a vocabulary size limit of 30K tokens.
A.2 T5
T5, or Text-to-Text Transfer Transformer, is a transformer-based architecture that casts all natural language processing (NLP) tasks into a text-to-text format. This means the model takes text as input and generates text as output, regardless of the specific task. At its core, T5 is an encoder-decoder model with the following key components:
  • Encoder: This component takes the input text and processes it into a sequence of hidden representations. It uses multiple transformer layers, each consisting of self-attention mechanisms and feedforward neural networks.
  • Decoder: This component generates the output text auto-regressively, conditioned on the encoder’s hidden representations. It also uses multiple transformer layers with self-attention and feedforward networks, as well as cross-attention mechanisms to attend to specific parts of the input sequence.
Datasets. T5 is pre-trained on a massive dataset called C4 (Colossal Clean Crawled Corpus), which contains around 750 GB of clean English text.
Tokenization. It uses Sentencepiece tokenizer on the cased corpus with a vocabulary size limit of 30K tokens.
A.3 GloVe
GloVe (Global Vectors for Word Representation) is a method for obtaining vector representations for words. Unlike context-based models like transformer-based LM models, GloVe leverages global word co-occurrence statistics across a corpus to learn word vectors. The GitHub repository [^8] https://github.com/stanfordnlp/GloVe provides an implementation of the GloVe model for learning word representations (word vectors or embeddings). We used the default embedding provided by the python API  [^9] Common Crawl (840B tokens, 2.2M vocab, cased, 300d vectors, 2.03 GB download), and a context window size of 10..
Tokenization. They used Stanford tokenizer  [^10] https://nlp.stanford.edu/software/tokenizer.shtml, a form of BPE tokenization scheme [47, 12] that constructs unigram counts from a corpus, and optionally thresholds the resulting vocabulary based on total vocabulary size (2.2M 1111112196016 cased tokens most frequent words in the case of GloVe embeddings) or minimum frequency count [40].
Appendix B Human-LM Alignments For Glove, T5
We performed our concept extraction algorithm on Glove, ALBERT, and T5 models. We observed that concept formation happens in all the models, however, the quality of the formed concepts is better in ALBERT model. Glove embeddings contain more than 200K tokens, in order to enable apple-to-apple comparison with ALBERT, we found the intersection of these tokens with the huggingface Albert-base-v2 token set, and, applied our concept extraction algorithm. We the formed clusters have weak correlation with the concepts formed in ALBERT model 121212if we assume Glove-ALBERT mapping exists between two clusters if more than half of their members are equivalent, %69 percent of the Glove clusters have a corresponding ALBERT cluster for K=6. (they similarities are stronger for concrete names/entities).
Tables 7 and 8 show the T5 location and name communities detected by our algorithm. High precision numbers for these cluster indicate clear concept formation. However,since the pretrained HuggingFace T5 uses cased token set for the pretraining, the number of tokens in associated with location and names are much smaller than T5 and Glove.
Tables 9 and 10 show the GloVe location and name communities detected by our algorithm. Although, it shows the categories are formed in GloVe embeddings as well, the numbers suggests the quality of the formed categories have less quality than both ALBERT and T5 counterpart.
Table 7: Location communities founds in T5 tokens. Note that most of the communities are member of 0_0_5, and 0_0_13 super communities.
Precision Support Cluster Name
USA 0.843 159 0_0_5_4
Britain/Ireland 0.843 118 0_0_5_5
Africa/South-America 0.672 64 0_0_13_2
Germany 0.619 21 0_0_13_0_0
Australia 0.842 19 0_0_5_10
France 0.867 15 0_0_13_0_2
Canada 1.000 13 0_0_5_14
Balkan 0.833 12 0_0_13_1_1
Indochina 0.818 11 0_0_13_3_1
Benelux 0.700 10 0_0_13_0_3
Canada 0.778 9 0_0_5_24
India 0.750 8 0_0_13_4_0
Romania 0.714 7 0_0_22_1_4
Central Europe 0.800 5 0_0_13_0_5
Israel/Palestine 0.800 5 0_0_13_5_3
Arab Countries 0.800 5 0_0_13_5_5
Nordic 0.750 4 0_0_13_1_6
Baltic 1.000 4 0_0_13_1_8
Table 8: Name communities founds in T5 tokens. Note that most of the communities are member of 0_0_5 super community.
Gender Overall Detected Country Support Cluster
Precision Country Precision Name
Male 0.942 United States 0.962 209 0_0_5_0
Female 0.888 United States 0.858 162 0_0_5_3
Last Name 0.904 United States 0.914 198 0_0_5_1
Male 0.837 France 0.612 49 0_0_5_7
Male 0.909 Peru 0.758 33 0_0_5_8
Mix 0.960 Germany 0.84 25 0_0_5_9
Male 0.846 Russian/Italy 0.615 13 0_0_5_15
Politician Names 1.0 N/A N/A 14 0_0_5_13
Table 9: Location communities founds in the subset Glove tokens that exists in Albert Vocab.
Country Precision Support Cluster
Names
US 0.674 522 0_4
UK 0.626 174 0_8
Europe 0.513 39 0_3_8
China 0.625 32 0_3_10
Italy 0.720 25 0_3_14
Philippines 0.556 18 0_3_15
Spain 0.722 18 0_3_15
Japan 0.688 16 0_3_16
France 0.857 14 0_3_9_0
Africa 0.750 12 0_3_3_1
Indochina 0.600 10 0_3_11_1
Netherlands 0.833 6 0_3_17_0
Table 10: Name communities founds in the subset Glove tokens that exists in Albert Vocab.
Country Overal Country Gender Support Cluster
Precision Precision Name
USA 0.835 0.555 female 575 0_2_1
UK 0.695 0.641 male 223 0_2_0_0
USA 0.737 0.337 male 95 0_2_6
Italy 0.942 0.692 male 52 0_2_4_1
Mexico/Colombia 0.889 0.711 male 45 0_2_4_2
Mexico/Peru 0.844 0.6 male 45 0_2_4_3
Austria 0.897 0.793 male 29 0_2_3_3
US/Nigeria 0.821 0.429 male 28 0_2_0_1_2
Russia 0.926 0.481 male 27 0_2_3_4
Saudi Arabia 0.8 0.84 male 25 0_1_6_1
Switzerland/Belgium 1 0.846 male 13 0_2_3_0_0
UAE 0.727 0.727 male 11 0_1_6_3
Saudi-Arabia 0.8 0.7 male 10 0_1_6_0_1
Germany 0.917 0.75 no-gender 12 0_2_3_2_2
UK/Canada 0.917 0.33 no-gender 12 0_2_0_1_5
UAE 0.6 0.467 no-gender 15 0_1_6_2
Appendix C Extracted Concept Hierarchies
Figure 4 shows the overall structure of hierarchical communities extracted by our proposed method. The cluster names were suggested by GPT-4 and corrected by a human supervisor. The green blocks are the ones that are discussed in this paper.
Figure 4: Visualization of the hierarchical Communities from ALBERT. The green blocks show the clusters that being evaluated and discussed in this paper.

### v3

@@
-Concept Formation and Alignment in Language Models: Bridging Statistical Patterns in Latent Space to Concept Taxonomy
+Aligned at the Start: Conceptual Groupings in LLM Embeddings
 Abstract
-This paper explores the concept formation and alignment within the realm of language models (LMs).
-We propose a mechanism for identifying concepts and their hierarchical organization within the semantic representations learned by various LMs, encompassing a spectrum from early models like Glove to the transformer-based language models like ALBERT and T5.
-Our approach leverages the inherent structure present in the semantic embeddings generated by these models to extract a taxonomy of concepts and their hierarchical relationships.
-This investigation sheds light on how LMs develop conceptual understanding and opens doors to further research to improve their ability to reason and leverage real-world knowledge.
-We further conducted experiments and observed the possibility of isolating these extracted conceptual representations from the reasoning modules of the transformer-based LMs.
-The observed concept formation along with the isolation of conceptual representations from the reasoning modules can enable targeted token engineering to open the door for potential applications in knowledge transfer, explainable AI, and the development of more modular and conceptually grounded language models.
+This paper shifts focus to the often-overlooked input embeddings – the initial representations fed into transformer blocks.
+Using fuzzy graph, k-nearest neighbor (k-NN), and community detection, we analyze embeddings from diverse LLMs, finding significant categorical community structure aligned with predefined concepts and categories aligned with humans.
+We observe these groupings exhibit within-cluster organization (such as hierarchies, topological ordering, etc.), hypothesizing a fundamental structure that precedes contextual processing.
+To further investigate the conceptual nature of these groupings, we explore cross-model alignments across different LLM categories within their input embeddings, observing a medium to high degree of alignment.
+Furthermore, provide evidence that manipulating these groupings can play a functional role in mitigating ethnicity bias in LLM tasks.
+Aligned at the Start: Conceptual Groupings in LLM Embeddings
+Mehrdad Khatir Virginia Tech khatir@vt.edu                        Sanchit Kabra Virginia Tech sanchit23@vt.edu                        Chandan K. Reddy Virginia Tech reddy@cs.vt.edu
 1 Introduction
-Language models are rapidly approaching human-level language abilities.
-However, these advancements raise important concerns - on the one hand, despite their immense potential, language models are susceptible to unexpected and potentially harmful behaviors such as hallucination [21], stereotyping [25], misinformation [37], and leakage of sensitive training data [29].
-On the other hand, the extent of their capabilities and the nature of their understanding is not yet known, thus capturing the attention of diverse groups, from policymakers to academics [16, 39, 3, 18].
-While these issues may seem different, they both stem from a common concern - the "black box" nature.
-The limited ability to understand how they reach their outputs raises fears about their true capabilities, potential biases, and unintended consequences.
-Researchers are actively studying the semantic representations learned by large language models (LLMs), aiming to understand the inner workings of language models.
-While initially used to study the learning dynamics of linguistic features and abstractions within LMs [50, 26], this research has expanded to explore how these models acquire and represent knowledge about the world.
-Some studies [38, 18, 1] have explored the ability of transformer-based models to learn representations of color, spatial, and temporal information.
-In contrast, other studies have raised concerns about the coherence and completeness of this knowledge [16].
-The work of [39] even suggests that language models may lack a true conceptual understanding of the real world, echoing the "stochastic parrot" critique [3].
-This debate stems from the fundamental question of whether language models merely capture statistical patterns in language or truly grasp the underlying concepts they represent.
-There are so many works on the semantic relations of the tokens but limited research directly addresses the question of whether these relations are grounded in a conceptual understanding of the real world, or merely reflect co-occurrence patterns in language data [2, 28].
-Objectives.
-To address the aforementioned gap, this paper investigates the process of concept formation in language models.
-We posit our work based on the argument that ‘Concepts’ are mental representations of categories, which group objects, events, or ideas based on shared features [45].
-More generic concepts that correspond to broader categories and share fewer defining features.
-This is how the hierarchical categories/concepts are formed.
-In an unsupervised world, when there is no label, categories translate into clusters/communities.
-Then, similar to [28], we argue that if neural models of language are to have any knowledge about concepts, it ought to be reflected in their token/word embeddings.
-Thus, communities/categories need to form in the semantic space in which their most relevant category names are often closely related to concepts that describe that category.
-Note that, while we often form concepts based on categories, not every category will necessarily lead to a clear and distinct concept in the real world.
-This phenomenon, known as concept alignment [43] will be the focus of our quantitative analysis.
-Methodology.
-Interpretability research traditionally use probing [36, 51].
-Probing often involves evaluating models on specific, often artificial tasks designed to isolate particular linguistic phenomena.
-Probing results might be sensitive to the specific design choices of the probe tasks, such as the choice of dataset, prompts, or metrics, making it difficult to extrapolate conclusions to the model’s overall capabilities in real-world scenarios.
-Thus, while effective for specific domains, they have inherently limited ability to provide broader insights into the extent of models’ underlying representations and reasoning capabilities beyond specific tasks they are tuned for [48] Therefore, we avoid using conventional probing, instead, we propose a concept extraction methodology based on graph creation (K-nearest neighbors), fuzzy weighting mechanism (borrowed from UMAP category theory [32]) and community/cluster detection algorithms to identify groups of highly interconnected data points, representing potential concepts.
-The performance of this approach hinges upon the existence of a meaningful, hierarchical structure within the input embedding layer.
-This method can potentially offer a deeper and more holistic understanding of the underlying knowledge encoded in these representations (as compared to probing).
-Models Under Investigation.
-Using our proposed method, we investigate the embedding spaces of GloVe, a traditional statistical language model [40]; ALBERT, a transformer-based encoder language model [23]; and T5, a transformer-based autoregressive language model [42] to comprehend if concept formation occurs in any of these models.
-Evaluation.
-In order to answer the question of "if concept formation happens in the semantic space," we evaluated the formed clusters against external references such as WordNet [34], Name Dataset [44], and country-state-city databases [11] for most of our evaluations.
-For clusters where open-source datasets were not available, we employed human annotators as an external reference.
-We observed that the process of concept formation can occur as early as the input embedding layer in transformer-based models.
-Considering that this observation hints at the existence of semantic memory in transformer-based models, we further investigate in Section 4.3 whether there is a separation between the identified semantic memory and the reasoning mechanisms.
-Specifically, we ask whether the model can still reason effectively with a modified semantic memory.
-This line of inquiry aims to deepen our understanding of the relationship between knowledge representation and reasoning capabilities within transformer-based models.
-Figure 1: Simplified steps on how external information is understood and retained.
-Upon understanding a newly encountered word/entity, it is typically stored in the semantic memory.
-The existence of semantic memory (on the left) allows the previously encountered words/entities to have a form of meaning even without requiring an external context.
-The scatter box on the right is the community (primarily associated with moving creatures) we extracted from the ALBERT model [23].
-2 Semantic Representation Space
-2.1 Concept formation, Representation Learning and Semantic Memory
-A simplified model of human understanding can be described as a multi-step process in which the incoming sensory information is integrated and turned into a brain-constructed interpretation of external objects/stimuli.
-This results in a mental construct (or percept) [15].
-Then, the newly formed mental construct gets compared, integrated, and associated with the existing knowledge already retained in the semantic memory.
-These connections help us to conceptualize and understand the new information based on what we already know (as shown in Figure 1).
-While diverse reasoning mechanisms exist, a common thread among them seems to be their reliance on and interaction with the semantic memory.
-This interaction likely leaves enduring traces within such a memory [5].
-The integration of newly encountered mental constructs within an existing cognitive framework is often guided by their relations to established internal constructs [13].
-This process facilitates the creation, refinement, or expansion of semantically similar clusters [27], implying a degree of inherent categorization [^1] This notion implies an interrelation between the recognition and categorization..
-Representation Learning and Concept Formation Representation learning is a fundamental aspect of language models, where the goal is to learn distributed representations (embeddings) for words or subword units that capture their semantic and syntactic relationships [4].
-Early models like word2vec [33] utilized shallow neural networks to learn these embeddings from large text corpora.
-More recent models, such as Large Language Models (LLMs) like GPT-3 [30], employ transformer architectures with self-attention mechanisms [53], enabling more accurate and dynamic representation learning.
-From the distributional hypothesis, models can form "concepts" by identifying patterns and relationships within data, particularly through recognizing the approximate invariance of shared features across different data instances [13].
-It should be noted that "concept formation" is a byproduct of the learning process, not a guaranteed outcome.
-Furthermore, the formed "concepts" might not always align with human-defined concepts.
-Thus, it is crucial to investigate and evaluate the nature of concept formation within models and their alignment with human understanding.
-2.2 Evaluation Methods
-The analysis, measurement, and interpretability of semantic representation learning in language models have been the subject of extensive research.
-Various methods have been proposed to evaluate how well these models capture semantic meaning.
-Intrinsic evaluation methods, assess semantic representations by measuring word embedding similarity, comparing them to human judgments of relatedness, (e.g.
-[33, 17, 52, 36, 51]).
-While these methods provide valuable insights, they often suffer from the limitations that probing techniques impose, and may not fully capture how well a model forms conceptual hierarchies or grasps the full spectrum of semantic relationships, as word embeddings typically represent individual words rather than higher-level categories [9].
-This limitation can hinder the evaluation of a model’s ability to understand broader semantic relationships and its capacity for abstract reasoning.
-Another approach is to use extrinsic evaluation methods, which measure the performance of language models on downstream tasks that rely on semantic understanding, such as sentiment analysis, question answering, and machine translation [54, 55].
-The performance on these tasks can indirectly indicate the quality of the learned semantic representations.
-The main challenge with using such extrinsic evaluation metrics to assess concept formation in the embedding space is that they do not directly measure the quality of the concepts themselves.
-Instead, they measure how well the model performs on specific tasks that rely on those concepts, which does not necessarily prove that the model has formed robust, human-like concepts.
-For instance, a model might accurately classify sentiment without truly grasping the nuances of emotions like sarcasm or irony [31].
-In addition to evaluation metrics, researchers have also employed visualization methods to interpret learned representations.
-These methods often involve projecting high-dimensional embeddings into lower-dimensional spaces for visualization [48, 46].
-However, several challenges arise with these approaches.
-Firstly, they offer indirect and subjective assessments of model understanding, lacking a quantitative basis for evaluation.
-Secondly, the dimensionality reduction process can lead to information loss [14, 56], potentially obscuring important nuances in the representations.
-Finally, different visualization techniques can produce conflicting results, making it difficult to reach definitive conclusions about the model’s true comprehension.
-To mitigate these shortcomings, we propose a concept extraction mechanism that identifies communities in higher-dimensional space.
-This approach enables both quantitative evaluation and mitigation of potential information loss incurred during dimensionality reduction.
+Large Language models (LLMs) are rapidly approaching human-level language abilities.
+However, these advancements raise important concerns - despite their immense potential, language models are susceptible to unexpected and potentially harmful behaviors such as hallucination 21, stereotyping 26, misinformation 36, and leakage of sensitive training data 29.
+Moreover, the extent of their capabilities and the nature of their understanding is not yet known, thus capturing the attention of diverse groups, from policymakers to academics 17, 39, 3, 19.
+While these issues may seem orthogonal, they stem from a common concern - the "black box" nature of large language models.
+The limited ability to understand how they reach their outputs raises doubts about their true capabilities, potential biases, and unintended consequences.
+A crucial step towards addressing these challenges lies in understanding how LLMs internally represent and manipulate semantic concepts, the building blocks of human language and thought.
+Unraveling these internal representations is essential for not only improving LLM performance and robustness, but also for building trustworthy machine learning systems.
+In the realm of contextual representations, some studies 38, 19, 2 have investigated the ability of transformer-based models to learn representations of color, spatial, and temporal information.
+These studies often rely on analyzing contextual embeddings, which are intermediate outputs of LLMs.
+Moreover, contextual embeddings are inherently tied to the specific input context, making them highly variable, less generalizable, and less controllable/manipulatable.
+In our work, we explore this critical area by investigating the emergence of conceptual structures within the input embedding space of LLMs.
+We aim to explore if these conceptual formations form independent of the context, and if such groupings are in alignment across LLMs, if these clusters exhibit internal organization, and if these patterns have functional implications for model behavior.
+Specifically, we examine:
+  • RQ1: Whether semantically related words and phrases are grouped together, forming identifiable conceptual clusters that are aligned with external world concepts and categories.
+  • RQ2: Do these groupings exhibit intra-cluster organization such as hierarchies, topological ordering, etc., thereby suggesting the formation of structured concepts.
+  • RQ3: Is there an inter-model alignment in semantic organization, across diverse transformer-based LLMs (Albert 23, T5 42, Llama3 1) irrespective of architecture, size, or pretraining data.
+  • RQ4: Whether these groupings play a functional role in LLMs. We test this by a case study that tries to mitigate ethnicity bias through cluster modification.
+To uncover conceptual structures within the input embedding space of LLMs, we employ fuzzy graph construction 32.
+Then, the fuzzy graph is analyzed using a community detection algorithm 6 to reveal conceptual groupings and their categorical organization.
+We use Louvain community detection, which is effective at revealing hierarchical community structures 6 in conjunction with multiple k choices (for k-nearest neighbors) in our approach.
+This systemic approach allows us to investigate both the existence and the hierarchical organization of conceptual clusters, directly addressing our research questions.
+Quantitative evaluation on external datasets (named entities 43, 12, numerical tokens, and social structures) demonstrates that token embeddings exhibit significant categorical community structure aligned with real-world concepts.
+The structure of this paper is as follows: We first establish necessary background on embeddings, semantic representations, and evaluation strategies (Section 2), followed by a description of our methodological approach (Section 3).
+We then present our core findings, starting with LLM-human alignment (Section 4), with a focus on within-cluster properties and hierarchical structure.
+Then, we explore LLM-LLM alignment within their input embedding space (Section 5).
+Section 6 demonstrates the practical implications of our work through embedding engineering and bias mitigation.
+Finally, we conclude in Section 7.
+2 Preliminaries
+2.1 Static, Contextual and Base Embeddings
+In this section, we clarify the distinctions between static, contextual, and base embeddings, which are crucial for understanding modern language models.
+  • Static embeddings (e.g., Word2Vec33, GloVe 40) are context-independent vector representations of words, meaning each word has a fixed embedding regardless of its surrounding text.
+This limits their ability to handle polysemy (words with multiple meanings).
+These embeddings are typically pre-trained on large corpora and can be used in various downstream tasks.
+Critically, for the context of this work, static embeddings are product of legacy LM models and not used as the input representations of modern transformer-based LLMs.
+They are not inputs to transformer blocks and thus have limited significance when it comes to applications such as mitigation techniques (e.g.
+embedding engineering) in LLMs.
+  • Contextual embeddings These are dynamic, context-dependent vectors.
+The embedding of a token is a function of its surrounding text, enabling the representation of nuanced meaning and resolving polysemy (e.g., Bert 11, Albert 23, GPT variants 42, 1).
+Different model layers capture different levels of contextual conditioning.
+However, this context-dependence limits direct interpretability and generalizability of individual token embeddings outside of specific contexts.
+  • Base Embeddings: The process of generating contextual embeddings starts with base embeddings, which provide the initial vector representation for each input token.
+These differ from static embeddings such as GloVe and Word2Vec from the following perspectives.
+(1) Generation: Base embeddings are learned parameters within the LLM, as compared to separately trained static embeddings.
+(2) Usage: They are the direct input to the transformer blocks, forming the basis upon which contextualized representations are built through the model’s subsequent layers.
+These embeddings are the focus of our study.
+2.2 Previous Works on Embedding Interpretability
+Previous research on interpretability in LLMs has primarily focused on analyzing either contextual embeddings (for modern LLMs) or static embeddings (for legacy language models which are not directly applicable to LLMs due to the architectural difference).
+In the realm of contextual representations, initial research focused on the learning dynamics of linguistic features within LLMs 50, 27, the scope has expanded to explore how these models acquire and represent knowledge about the world.
+Some studies 38, 19, 2 have investigated the ability of transformer-based models to learn representations of color, spatial, and temporal information.
+These studies often rely on analyzing contextual embeddings, which are intermediate outputs of LLMs.
+Moreover, contextual embeddings are inherently tied to the specific input context, making them highly variable and less generalizable.
+The closest work to our knowledge that focuses on context agnostic embeddings within the modern LLMs is 7 where the authors propose a method to create context agnostic word embeddings from contextualized word representations using (sub)-word pooling as well as context combination techniques, and tested on semantic similarity datasets.
+Furthermore, 25 proposes a method for creating context-agnostic word representations by averaging the contextual embeddings derived from BERT, given a set of inferences on a masked target token within a corpus.
+In this methodology, the context is seen as a form of Gaussian noise that can be averaged out and hence produces a context-agnostic semantic representation.
+They observed such embedding represent richer semantic information than static word embedding counterparts (Word2Vec and GloVe) in intrinsic evaluation tasks.
+However, these works lack rigorous and extensive analysis of base embeddings of the LLMs to explores the intrinsic semantic organization within input embeddings of LLMs, possbile conceptual groupings and their alignments.
 3 Concept Extraction
-As discussed in Section 2.1, identifying clusters or communities within the semantic representation space is a crucial initial step in evaluating concept formation.
-We prioritize community detection among clustering methods for the following reasons:
-  • Alignment with Semantic Network Structure: The underlying semantic representation, as detailed in Section 2.1, exhibits a network-like structure.
-Community detection algorithms are specifically designed to uncover such inherent community structures within networks, making them better suited to our task.
-  • Unknown Number of Clusters: Traditional clustering algorithms are often sensitive to the choice of the number of clusters or other clustering parameters [57].
-In our case, we lack this prior knowledge and only assume that semantically similar words are closer in the representation space.
-Community detection methods do not necessitate a predetermined number of clusters.
-  • High Dimensionality of Data: The high dimensionality of the semantic representation space poses challenges for many clustering algorithms [57].
-Community detection methods on the other hand, naturally circumvent this issue by transforming the high-dimensional space into a graph , where relationships between data points become more readily apparent.
-This approach allows for a more effective and accurate identification of semantic clusters, reflecting the natural structure of the semantic representation space and facilitating a deeper understanding of concept formation.
+To investigate RQ1 and RQ2, we study the Human-LLM alignment of the input representations.
+To this end, we develop a method to first extract possible formed concepts within that space (refer to appendix A.1 for the discussion of conceptual groupings and semantic memory), and then to evaluate them against external datasets.
+Our methodology consists of building the semantic graph, then using community detection to extract possible conceptual groupings.
+[^1] Note that we favor community detection over traditional clustering for identifying the conceptual clusters due to (1) its ability to align with the network-like structure of semantic representations, (2) its independence from the need for a predetermined number of clusters, and (3) its effectiveness in managing high-dimensional data by transforming it into a graph.
+Note that differentiable embedding functions guarantee a smooth, optimizable embedding space, but do not ensure a uniform distribution of concept instances.
+This unevenness, arising from factors like varying concept complexity and instance frequency, implies that while conceptual groupings may be locally uniform, the overall embedding space can be unevenly distributed.
+Therefore, effectively identifying these groupings requires mitigating this unevenness, which we address using a UMAP-based fuzzy graph construction.
+Data: All tokens in the input embeddings.
+Result: A set of hierarchical communities.
+1 Create a community list. The initial community is the entire space;
+2 for k=[different neighbor sizes] do
+3       for all communities do
+4             - Generate knn graph from the input embedding weights;
+5             - Compute the edge weights of the graph using fuzzy simplex;
+6             - Apply Louvain community detection;
+7             - Add the identified communities to the list;
+8
+9
+Algorithm 1 Concept Extraction
 3.1 Graph Construction
-While categories themselves are not strict clusters, they may exhibit internal structure or fuzzy partitions based on shared features or relationships.
-Considering our discussion in the previous section, we seek a representation space where the structure of these categories is explicitly reflected in the geometry of the manifold.
-This requirement arises from the desire to directly manipulate and modify these clusters through geometric operations on the representation, opening up possibilities for Token Engineering within the language models.
-Inspired by the effectiveness of the UMAP dimensionality reduction [32], we use the Nearest neighbor descent to find the $K$ nearest neighbors for every token embedding in the embedding space and then use the following equations as the weight function of the edge between $x_{i}$ and $x_{j}$ nodes in the $K$-NN graph [32]:
+The first step in the Uniform Manifold Approximation and Projection (UMAP) algorithm is to approximate the manifold by constructing a fuzzy topological representation of the embeddings using nearest neighbor descent 32.
+Inspired by that, we use the same nearest neighbor descent method to find the $K$ nearest neighbors for every token embedding in the embedding space and then use the same equations used in UMAP’s fuzzy graph construction to define the weight function of the edge between $x_{i}$ and $x_{j}$ nodes in the $K$-NN graph 32:
 $\omega(x_{i},x_{j})=exp\left(\frac{-\max(0,~{}d(x_{i},x_{j})-\rho_{i})}{\sigma_{i}}\right)$ (1)
@@
 $\sum_{j=1}^{k}\omega(x_{i},x_{j})=\log_{2}(k).$ (3)
-Building upon our theoretical arguments, the invariant representations (if they exist), should form fuzzy partitions that can be detected by the community detection algorithms.
-Thus, for the third step, we perform community detection.
+Building upon our theoretical arguments, the conceptual/categorical representations (if they exist), should form fuzzy partitions that can be detected by the community detection algorithms.
+Note that since UMAP dimensionality reduction process can lead to information loss 15, 56, potentially obscuring important nuances in the representations, we perform community detection in the high-dimensional space.
 3.2 Louvain Community Detection
@@
 The modularity of a partition is a scalar value between -1 and 1 that measures how much more densely connected the nodes within a community are compared to how connected they would be in a random network.
-[6].
+6.
 For a weighted graph, modularity is defined as:
@@
 where $A_{ij}$ represents the weight of the edge between $i$ and j; $k_{i}$ is the sum of the weights of the edges attached to vertex $i$; $m$ is the sum of all of the edge weights in the graph; the $\delta$-function $\delta(c_{i},c_{i})$ is 1 if $u=v$ and 0 otherwise; $c_{i}$ is the community to which the nodes $i$ belongs to.
-Then, it aggregates the communities to identify hierarchical structures.
+Then, it aggregates the communities to identify possible hierarchical structures.
 In this phase, each community is considered as a single node and the links between the new nodes are calculated as the sum of the weight of the links between nodes in the corresponding two communities.
-More details are given in Algorithm 1.
+More details are given in appendix C, algorithm 2.
+3.3 Concept Extraction Algorithm
+For our concept extraction algorithm, as the first step, we create and weight the adjacency graph using $K$-NN, UMAP-based weighing formula (mentioned in section 3.1), and then use Louvain algorithm.
+Algorithm 1 describes the concept extraction process (see appendix E for details on the algorithm methodology and considerations).
+We configured our algorithms to create $k$-NN graph iteratively for different values of $k$.
+This enables us to observe the communities/concepts formation at various granularities.
+Table 1 shows the number of identified clusters for $k=[100,75,50,25,12,6]$.
+Table 1: Number of communities with different granularities of nearest neighbors for Albert, T5, Glove, and Llama3. For Glove, we only used the subset of GloVe that present in Albert vocabulary
+$K$ settings Vocab
+Models 100 75 50 25 12 6 Size
+T5 1 115 1203 4551 8137 9407 32000
+Albert 8 133 1058 4442 7718 8626 30000
+GloVe 9 207 1157 3521 6237 7200 25869
+Llama-3 7 23 844 6044 18644 32535 128256
+Hierarchy Formation.
+When examining the broader perspective (i.e., $k$=100), the model primarily found groups of named entities (names of people and places), adverbs, sub-words, some number symbols, and etc.
+(appendix J, Figure 5 shows the overall concept hierarchies of the Albert vocab).
+Zooming in further (e.g., $k$=75), these communities revealed more specific clusters that are relatable to the real world.
+For instance, within named entities, clusters formed for personal vs.
+location names, even further pinpointing locations by country.
+As the granularity level increases (approaching a smaller $k$ value), clusters exhibit a stronger association with word forms.
+Figure 1: The identified name and location communities for different k granularity for Albert model.
+At the left-side the average precision score for the extracted graph within each granularity is given.
+For the more detailed tables, see H (Note that the results for other models are also available in appendix I).
+Figure 2: Visualization of the identified name and location communities of size larger than 10 entities. UMAP projection along with Seaborn 57 is used for the visualization.
+4 Evaluation: Alignment with External Knowledge
+We extracted several meaningful conceptual communities, ranging from symbo-numerical groups, to concrete objects and Named-Entities (such as plant groups, animals, car brands, names, and locations), to more abstract groupings such as colors, social roles and structures, currencies, etc.
+(more details in appendix J).
+To evaluate RQ1 and RQ2, we selected named entities as well as numbers, primarily because the existence of external evaluation datasets for NE as well as the well defined nature of numbers allow us to examine (1) the alignment with external world knowledge, and (2) to observe within group properties of the conceptual communities.
+To assess the consistency of conceptual groupings across different LLMs, we applied our method to the embedding spaces of GloVe 40, Albert 23, T5 42, and Llama-3 1 (See Appendix F).
+For brevity, we present the overview of the results for Albert in the main section (for more detailed table for the all aforementioned models please refer to Appendix H and I)
+4.1 Named Entities
+To evaluate the quality and the alignment of the extracted conceptual grouping with the external world, we used we leveraged the name-dataset 43 and the country-state-city database 12.
+These datasets provide comprehensive information on names (including gender and country-specific popularity) and locations (including hierarchical relationships between countries, states, and cities), enabling a rigorous assessment of the conceptual groupings (see appendix H for more details).
+The granularity of observed clusters was influenced by both the number of constituent samples and conceptual attributes.
+For instance, English/American named entities and locations tended to form distinct clusters at higher levels of granularity (k values) due to their high frequency within the vocabulary.
+However, upon further increasing granularity (higher k values), hierarchical structure emerged within these clusters.
+The conceptual attributes contributing to this hierarchical organization within named entities included (1) entity type (human name or location), (2) name part (first or last), gender, and (3) regional/national origin.
+Note that since there’s no one-to-one mapping between names/locations in LLM vocabulary and external datasets (e.g., a name may appear in multiple countries, the external dataset is also a superset), recall is less relevant.
+We prioritize precision to evaluate accuracy, as it better reflects our ability to identify correct matches.
+We observe high precision score across the identifies name entity cluster.
+Figure, 1 shows the main groupings for Albert Model.
+We got average precision score of 0.81 for K=75 and 0.76 for K=50/25.
+Appendix H.1 contains the detailed tables for different models.
+Furthermore, we observed a degree of geographic ordering within the identified communities.
+As illustrated in Figure 2, there appears to be a general trend from east to west as we move across the communities from bottom-left to top-right.
+The leftmost communities are predominantly associated with Japanese locations and names, while those on the rightmost side are primarily linked to Europe and the United States.
+This suggests that the model’s internal representations in the input embedding layer may inherently capture geographical relationships.
+Table 2: Topological ordering score of different communities of the numbers. Years refer to the cluster of numbers between 1816 - 2021. Support refers to the number of samples within the cluster.
+Category k-1 k-3 k-5 Support
+0-100 0.68 0.89 0.92 100
+100-262 0.75 0.94 1.0 115
+263-300 0.46 0.78 0.89 37
+300-400 0.42 0.71 0.84 82
+400+ 0.53 0.75 0.83 110
+Years 0.86 0.96 0.98 203
+4.2 Symbols-Numbers
+At a high level, symbols and numbers are distinctly separated from other tokens in the embedding space.
+Within their own domain they form communities according to (i) years, (ii) integer values, (iii) tokens indicating monetary values (e.g., \$1), (iv) ratio/time (e.g., 2:1, 3:30), (v) fractions (e.g., ’/8’), (vi) large values (e.g., ‘100,000’), and (vii) percentages (e.g., $42\%$).
+In these communities, integers create sub-communities based on their hundreds.
+To align with the human notion of numbers, it is essential to identify a topological ordering within the numerical embeddings.
+As the embedding manifold may not conform to a Euclidean structure 24, 9, 8, conventional distance-based order measures were inapplicable.
+Lemma 5 formulates the local ordering of embeddings within a embedding manifold.
+Lemma 1.
+Local Ordering on a Manifold: Let M be a manifold and let d(a, b) denote the distance between points a and b on M.
+For a given positive integer $k$, we say that a point x on M is locally ordered if and only if:
+$x\in top_{k}(x+1)\cap top_{k}(x-1)$ (5)
+where $top_{k}(a)$ denotes the set of k-nearest neighbors of point a on M.
+This localized approach captures the intuitive concept of topological ordering within a small neighborhood, rather than relying on a global ranking.
+Topological Ordering Score is defined as:
+$S=\frac{1}{n}\sum\limits_{i=0}^{n}{f_{k}(x_{i})}$ (6)
+where $n$ indicates the number of embeddings in the given cluster, $k$ is an integer that controls the strictness of ordering, and $f_{k}(x_{i})$ is boolean function that returns 1 if $x_{i}$ hold lemma 5 condition.
+Table 2 presents the topological ordering scores of the communities identified by our algorithm.
+The "k-1" score represents strict topological ordering, where $x+1$ lies exactly between $x$ and $x+2$ in the embedding space.
+The "k-3" and "k-5" scores measure relaxed topological ordering, where $x+1$ falls within the 3 or 5 nearest neighbors of both $x$ and $x+2$, respectively.
+As the results indicate, all detected communities exhibit a high degree of local topological ordering, regardless of the chosen level of strictness.
+This finding is significant because it suggests that LLMs may possess not only the ability to categorize heterogeneous input entities but also the capacity to construct meaning within smaller, internally consistent structures (internally homogeneous sub-structures).
+This implies the potential for LLMs to move beyond simply classifying information to actively interpreting and generating meaning within specific clusters.
+5 LLM-LLM alignment
+Investigating LLM-LLM alignment in the input embedding layer is crucial to answer RQ3, because it reveals how well different LLMs represent the same concepts in their initial processing stages.
+This analysis provides valuable insights into the influence of model architecture, size, and training regimes on the formation of language representations.
+To quantify this alignment, we defined an alignment score that specifically captures the discrete overlaps of nearest neighbors in embedding spaces.
+This score offers a precise assessment of how similarly the LLMs represent shared tokens, enabling a more nuanced understanding of their representational alignment.
+We define Alignment Score as:
+$S=\frac{1}{n}\sum\limits_{j=0}^{j=n}\frac{|top\_k({LLM}^{1}_{j})\cap top\_k({LLM}^{2}_{j})|}{k}$ (7)
+where ${LLM}^{1}$ and ${LLM}^{2}$ denote the two LLMs under investigation, $n$ represents the number of shared tokens between the LLM pair, and $top\_k({LLM}^{1}_{j})$ and $top\_k({LLM}^{2}_{j})$ represent the sets of the top k nearest neighbors of the $j$-th token in ${LLM}^{1}$ and ${LLM}^{2}$ respectively.
+The alignment score for two randomly chosen embeddings with the same vocabulary is 0 (assuming $k<<n$).
+We calculated this alignment score across a range of LLMs, encompassing diverse architectures and training objectives.
+we first identify the set of tokens shared across the vocabularies of a given pair of LLMs.
+For each shared token, we compute its k nearest neighbors ($k=3,5,10,50$) in the embedding space of both models.
+We use pairwise cosine similarity measure to find top-k tokens within an LLM.
+This selection included Albert (encoder-based), T5 (encoder-decoder), Llama 70B (decoder-only) to examine the impact of high-level model architecture.
+To investigate the effect of model size, we included T5 models of different scales: T5-small, T5-base, T5-large, T5-3B, and T5-11B.
+Table 3: Pair-wise alignment score for various LLM sizes, and training regimes. higher alignment score is better.
+Model-1 Model-2 k=3 k=5 k=10 k=50 Support
+Albert-XXL Albert-base 0.501 0.490 0.472 0.438 30000
+Albert-XXL Albert-L 0.56 0.54 0.52 0.48 30000
+Albert-XXL Albert-XL 0.60 0.58 0.55 0.51 30000
+T5-11B T5-small 0.61 0.577 0.52 0.38 32100
+T5-11B T5-base 0.68 0.66 0.60 0.51 32100
+T5-11B T5-large 0.74 0.71 0.66 0.61 32100
+T5-11B T5-3B 0.79 0.76 0.71 0.63 32100
+Llama3-70B Llama3-1B 0.57 0.55 0.50 0.31 128256
+Llama3-70B Llama3-3B 0.56 0.53 0.48 0.27 128256
+Llama3-70B Llama2-70B 0.65 0.61 0.56 0.47 22430
+Table 4: Cross architecture alignment score for various LLM architectures, sizes, and training regimes.
+higher alignment score is better.
+Support column indicates the number of shared tokens between the LLM pair.
+Model-1 Model-2 k=3 k=5 k=10 k=50 Support
+Albert-XXL Llama3-3B 0.55 0.52 0.49 0.44 18640
+Albert-XXL T5-3B 0.63 0.61 0.59 0.54 12307
+Llama3-3B T5-3B 0.66 0.62 0.58 0.53 20603
+Llama3-70B T5-11B 0.64 0.61 0.56 0.47 20603
+Analyzing results from tables 3 and 4, we observe:
+  • Table 3 shows that model size is a contributing factor in achieving higher alignment scores when comparing within the same architecture.
+Since larger models also demonstrate better performance on benchmark tasks, we can infer that the quality of the concepts formed in the input embedding layer is positively correlated with model size, assuming architecture and training regimes are held constant.
+  • Both tables 3 and 4 show that alignment scores generally decrease as the value of k increases.
+This trend suggests that while models share a core understanding of semantic similarity at smaller k values, they diverge when considering more generic concepts at larger k values.
+The alignment drop is larger in decoder-only models, likely due to their unidirectional context  [^2] This is consistent with human studies showing that readers’ eye movements are bidirectional: forward to absorb new information, and backward to resolve comprehension issues or correct errors 48..
+  • Interestingly, the alignment scores between Llama2-70B and Llama3-70B are comparable to those between Llama3-70B and T5-11B (refer last rows of table 3 and table 4).
+There are minimal architectural differences between Llama2 and Llama3-70B (primarily in the context window size).
+This implies that factors such as training regimes (e.g., dataset size and composition) and context window size are as influential as model architecture in achieving strong alignment.
+  • Table 4 shows the cross-architecture alignment scores for different LLMs, comparing models of similar sizes but varying architectures.
+Despite architectural differences, models of similar sizes exhibit moderate to high alignment scores (mostly above 0.5), suggesting a consistent semantic organization in LLMs.
+Overall, we observed moderate to high alignment across LLMs, regardless of their size, architecture, or pretraining, indicating that the findings in Section 4 may generalize to other LLMs.
+6 Bias Mitigation: Case Study of Cluster Modification
+To investigate RQ4 and demonstrate the practical impact of our findings, we conduct a case study on modifying conceptual clusters to mitigate ethnicity bias.
+A key challenge here is balancing bias reduction with the preservation of the model’s overall utility and linguistic integrity.
+Our approach focuses on embedding engineering, targeting clusters of tokens associated with stereotype-prone identities, such as proper nouns from the Indian subcontinent (e.g., Indian and Pakistani human names).
+We hypothesize that by modifying these token embeddings, we can disrupt learned biases without sacrificing the tokens’ semantic roles within the language model.
+For example, we aim to investigate if the token “Sharma”, a member of the Indian proper nouns cluster, retains its identity as a proper noun after token engineering.
+To evaluate fairness score, we use Bias Benchmark for Question Answering (BBQ) dataset 37, the fairness evaluation methodology is based on the WinoBias evaluation59.
+To assess the linguistic preservation, we chose the part-of-speech (POS) tagging task as a proxy indicator on CoNLL-2003 dataset 52 as well as a subset of Wikimedia dataset 58 (Further evaluation on the robustness of embedding engineering on GLUE, SuperGLUE, and SQUAD 55, 54 benchmarks is provided in Appendix D to provide a more comprehensive assessment of general language model quality after the token manipulations).
+Our approach begins by selecting a cluster of token embeddings from the original model.
+For this experiment, we used the communities associated with Indian and Pakistani human names that was identified by our concept extraction algorithm.
+Then, we calculated the mean and standard deviation of the joint cluster.
+Finally we modified these clusters to form a single cluster with Gaussian distribution of the joint cluster (i.e.
+samples from the $Gaussian(\mu,0.7*\sigma)$).
+Then, we finetuned the model for POS tagging task for CoNLL-2003 dataset fro 5 epochs (see Appendix B for more details).
+For the POS tagging task, we utilize the CoNLL-2003 dataset 52, a widely-used benchmark for named entity recognition and POS tagging.
+Each model is fine-tuned using a standard supervised fine-tuning approach.
+As shown in Table 5, all token-engineered models exhibit over 90% token overlap with their base counterparts.
+This high overlap (like the >90% reported) suggests that even though the embeddings have been modified, they still largely represent the same underlying POS tags.
+Furthermore, for those POS tags where the base and modified models show disagreement, we still observe the same overall quality.
+This means that even where the model does change its POS tag assignment, the new assignment is just as likely to be correct as the original.
+At the same time, we are able to mitigate bias ranging from 44% for the Albert-xxl-large and T5-3b models to 52% for the T5-11b model.
+These highlight the potential of token engineering for bias mitigation while keeping the general purpose utility of the model intact.
+Table 5: Fairness and POS Tagging Performance of Various Models. Fairness: higher score implies less bias. POS: higher score implies better accuracy.
+Fairness POS Tagging
+Model Base Ours Base Our Overlap
+Score Score Acc Acc %
+Albert-base 0.26 0.74 0.91 0.91 0.90
+Albert-xxl 0.28 0.72 0.93 0.91 0.92
+T5-3b 0.28 0.72 0.92 0.91 0.94
+T5-11b 0.24 0.76 0.94 0.93 0.94
+7 Conclusion
+In this paper, we propose a modular concept extraction mechanism that uncovers the emergence of distinct conceptual communities within the entire input embedding space.
+Using our methodology, we observe that LLMs form organized conceptual structures within their input embedding spaces.
+We demonstrate that the input embeddings of LLMs form categorical semantic structures that align with external world representations.
+We quantitatively analyze several properties of these structures, with a particular focus on categorical structures related to named entities.
+Additionally, we observe that numerical structures within the input embedding layer align with human notion of numerical values, including a topological ordering of numbers.
+We also discussed that LLMs inherently exhibit a degree of alignment with one another, suggesting the potential to extend the observed human-LLM alignment to other models.
+This study opens new avenues for further exploration and intervention in LLMs, especially within the realm of embedding engineering in several key areas, including bias detection and mitigation.
+8 Limitations and Risks
+Limitations: The model forms the conceptual communities that are meaningful but its priority is not exactly the same as that of humans.
+The model vocabulary is a contributing factor to the way the model prioritizes the formation of conceptual clusters in its embedding layer.
+For example, the number of English names is much higher than the other languages and this has caused the model to form high-level communities (e.g., $k$=75) specified for names vs.
+less frequent names/locations a high-level community contains the combination of regions personal and location names.
+This limits our method to associate the KNN resolution with the abstraction level of the extracted concepts/categories.
+Risks: This work provides detailed information about (1) the formed clusters/concept in the input embedding layer, and (2) the separation of memory from reasoning in Albert.
+As the methodology can also be applied to other models, it can potentially facilitate more advanced adversarial attacks and content manipulation in LLMs.
+Appendix A Semantic Representation Space
+A.1 conceptual groupings, Representation Learning and Semantic Memory
+A simplified model of human understanding can be described as a multi-step process in which the incoming sensory information is integrated and turned into a brain-constructed interpretation of external objects/stimuli.
+This results in a mental construct (or percept) 16.
+Then, the newly formed mental construct gets compared, integrated, and associated with the existing knowledge already retained in the semantic memory.
+These connections help us to conceptualize and understand the new information based on what we already know (as shown in Figure 3).
+While diverse reasoning mechanisms exist, a common thread among them seems to be their reliance on and interaction with the semantic memory.
+This interaction likely leaves enduring traces within such a memory 5.
+The integration of newly encountered mental constructs within an existing cognitive framework is often guided by their relations to established internal constructs 14.
+This process facilitates the creation, refinement, or expansion of semantically similar clusters 28, implying a degree of inherent categorization [^3] This notion implies an interrelation between the recognition and categorization..
+Representation Learning and Conceptual roupings) Representation learning is a fundamental aspect of language models, where the goal is to learn distributed representations (embeddings) for words or subword units that capture their semantic and syntactic relationships 4.
+Early models like word2vec 33 utilized shallow neural networks to learn these embeddings from large text corpora.
+More recent models, such as Large Language Models (LLMs) like GPT-3 30, employ transformer architectures with self-attention mechanisms 53, enabling more accurate and dynamic representation learning.
+From the distributional hypothesis, models can form "concepts" by identifying patterns and relationships within data, particularly through recognizing the approximate invariance of shared features across different data instances 14.
+It should be noted that "conceptual groupings" is a byproduct of the learning process, not a guaranteed outcome.
+Furthermore, the formed "concepts" might not always align with human-defined concepts.
+Thus, it is crucial to investigate and evaluate the nature of conceptual groupings within models and their alignment with human understanding.
+A.2 Evaluation Methods
+The analysis, measurement, and interpretability of semantic representation learning in language models have been the subject of extensive research.
+Various methods have been proposed to evaluate how well these models capture semantic meaning.
+Intrinsic evaluation methods, assess semantic representations by measuring word embedding similarity, comparing them to human judgments of relatedness, (e.g.
+33, 18, 51, 35, 49).
+While these methods provide valuable insights, they often suffer from the limitations that probing techniques impose, and may not fully capture how well a model forms conceptual hierarchies or grasps the full spectrum of semantic relationships, as word embeddings typically represent individual words rather than higher-level categories 10.
+This limitation can hinder the evaluation of a model’s ability to understand broader semantic relationships and its capacity for abstract reasoning.
+Another approach is to use extrinsic evaluation methods, which measure the performance of language models on downstream tasks that rely on semantic understanding, such as sentiment analysis, question answering, and machine translation 55, 54.
+The performance on these tasks can indirectly indicate the quality of the learned semantic representations.
+The main challenge with using such extrinsic evaluation metrics to assess conceptual groupings in the embedding space is that they do not directly measure the quality of the concepts themselves.
+Instead, they measure how well the model performs on specific tasks that rely on those concepts, which does not necessarily prove that the model has formed robust, human-like concepts.
+For instance, a model might accurately classify sentiment without truly grasping the nuances of emotions like sarcasm or irony 31.
+In addition to evaluation metrics, researchers have also employed visualization methods to interpret learned representations.
+These methods often involve projecting high-dimensional embeddings into lower-dimensional spaces for visualization 46, 44.
+However, several challenges arise with these approaches.
+Firstly, they offer indirect and subjective assessments of model understanding, lacking a quantitative basis for evaluation.
+Secondly, the dimensionality reduction process can lead to information loss 15, 56, potentially obscuring important nuances in the representations.
+Finally, different visualization techniques can produce conflicting results, making it difficult to reach definitive conclusions about the model’s true comprehension.
+To mitigate these shortcomings, we propose a concept extraction mechanism that identifies communities in higher-dimensional space.
+This approach enables both quantitative evaluation and mitigation of potential information loss incurred during dimensionality reduction.
+Appendix B Bias Mitigation details
+We evaluated the models on subsets of the Wikimedia dataset 58, where we sample only those sentences that contain the tokens from our token-engineered cluster.
+For the fairness task, we adopt the Bias Benchmark for Question Answering (BBQ) dataset 37, focusing on the nationality split.
+We modify the data set for masked language modeling (MLM) by replacing the interrogative questions with a token ’[MASK]’.
+To ensure the quality of evaluation, each sample is manually checked for grammatical correctness following this transformation.
+For evaluating fairness, we compare each base model with its corresponding token-engineered model that utilizes Gaussian-sampled embeddings.
+We first filter out the evaluation samples that pass the fairness test on the base model, thus isolating only the problematic cases.
+Our evaluation of fairness is inspired by the evaluation metrics used for the winobias dataset 59.For the remaining biased samples, we calculate which model (base or token-engineered) is more likely to generate a biased output by examining their output logit probability.
+Appendix C Lovain Algorithm
 Data: The initial input is a weighted network of all the nodes exist in the entire space.
@@
 15       Reduce each community to a single node;
-16       go to 1
+16       go to 2
 Algorithm 2 Louvain
-3.3 Concept Extraction Algorithm
-for our concept extraction algorithm, as the first step, we create and weight the adjacency graph using $K$-NN, UMAP-based weighing formula (mentioned in section 3.1), and then use Louvain [6] and label propagation [58] community detection algorithms.
-There are multiple viable strategies to extract hierarchical communities in our methodology; the first one is to use algorithms such as Louvain that inherently generate hierarchical communities [6].
+Evaluation Metrics.
+conceptual groupings involves the creation of abstract internal representations through the clustering of inputs with shared features.
+While numerous intrinsic clustering metrics exist to assess cluster formation quality, their application to our use case is limited by two factors:
+  • The formed clusters are not situated within Euclidean space 24, 9, 8, rendering geometric properties such as cluster distances inadequate indicators of concept well-formedness and distinctiveness.
+  • Concepts inherently possess a degree of vagueness 20, thus metrics like compactness or separation do not reliably reflect the quality of formed concepts.
+Appendix D Knowledge-Reasoning Separation
+Thus far, we have demonstrated that the LM constructs a knowledge base (mental representation) directly within its input embedding layer.
+Furthermore, we have established a degree of human-LM alignment in both hierarchical structure and semantic meaning.
+Table 6: Huggingface Albert base model on GLUE, SuperGLUE, and SQUAD tasks.
+For the baseline, the model was finetuned without altering the embeddings.
+For the mid-point, the embedding layer entries are assigned the mid-point embedding of their associated community.
+We used 41 repo for benchmarking.
+mid-point GLUE
+Method/Tasks mnli mrpc qnli qqp rte sst stsb wnli
+baseline 0.827 0.841 0.902 0.858 0.765 0.915 0.872 0.549
+mid-point 0.849 0.865 0.910 0.875 0.783 0.922 0.890 0.563
+mid-point SuperGLUE SQUAD SQUAD
+Method/Tasks boolq cb copa multirc record wic wsc v1 (f1) v2 (f1)
+baseline 0.622 0.512 0.59 0.350 0.586 0.595 0.528 83.72 70.9
+mid-point 0.621 0.478 0.55 0.372 0.588 0.626 0.634 84.4 74.9
+Now, we are interested to see the extent to which the knowledge learned during pretraining is modular and separable from the reasoning mechanisms employed by the language model in downstream tasks.
+For example, can the knowledge learned during the pretraining phase be selectively removed or modified without significantly impacting the model’s performance on finetuning?
+The modularity can also impact the effectiveness of Language Model Inversion 34, 47 techniques, which aim to extract private information such as names or other sensitive details learned during the pre-training444 For example, if knowledge is found to be highly modular, it may be possible to develop targeted interventions that obscure or remove specific sensitive information without significantly impacting the model’s overall performance on downstream tasks..
+To investigate this, we selected GLUE, SuperGLUE, and SQUAD benchmarks as downstream tasks to assess language model performance.
+We then systematically removed within-community information by calculating and assigning the embedding space center (mid-point value) of each community to all its members.
+For example, in a community of names like "James," "John," and "Alex," all members would share the same embedding.
+Table 6 shows the results on major LM benchmarks GLUE, SuperGLUE, and SQUAD for the Albert base model.
+Although our experiment does not prove the separation of knowledge and reasoning, it indicates that at least the granular information acquired during the pretraining is not required for the model’s performance on the aforementioned LM benchmarks.
+This is significant because it opens the door for embedding engineering of private or harmful information that is learned during the pretraining.
+Appendix E Hierarchical Community Extraction: Methodology and Considerations
+There are multiple viable strategies to extract hierarchical communities in our methodology; the first one is to use algorithms such as Louvain that inherently generate hierarchical communities 6.
 However, the summation of the weights as a new weight in the community aggregation phase of Louvain algorithm skews the weighted graph in favor of merging smaller communities in the next phase.
@@
 Thus, for our concept extraction algorithm, we use 1-2 community aggregation and rather use KNN iteratively with different granularity for extracting hierarchical concepts (more details are described in Algorithm 1).
-Note that we only use well-established methods such as $K$-NN, UMAP-based weighing formula, as well as Louvain [6] and label propagation [58] community detection algorithms to capitalize on the established generalizability of these algorithms.
+Note that we only use well-established methods such as $K$-NN, UMAP-based weighing formula, as well as Louvain 6 and label propagation 60 community detection algorithms to capitalize on the established generalizability of these algorithms.
 Although, as we show in the next sections that our method produces amazingly good categories, it should be noted that we intend to focus our analysis on “if the language model forms concepts" rather than creating the most optimal concept extraction mechanism.
@@
 Notably, our concept extraction is algorithm-agnostic; alternatives could be readily employed.
-Data: All tokens in the input embeddings.
-Result: A set of hierarchical communities.
-1 Create a community list. The initial community is the entire space;
-2 for k=[different neighbor sizes] do
-3       for all communities do
-4             (1) Generate KNN graph of the community using the input embedding weights;
-5             (2) Compute the edge weights of the graph using fuzzy simplex;
-6             (3) Apply Louvain/Label-Propagation community detection;
-7             (4) Add the identified communities to the list;
-8
-9
-Algorithm 2 Concept Extraction
-3.4 Models Under Investigation
-Using our proposed method, we investigated the embedding spaces of GloVe, ALBERT, and T5 to determine if concept formation occurs consistently across these diverse language models.
-We chose this spectrum of models to: (i) assess if our method for identifying concept formation is applicable across a wide range of architectures and training data, and (2) investigate whether the concept formation is a fundamental property of language representation, regardless of the specific model architecture or training data.
-We hypothesize that the language model with the least embedding layer dimension is more appropriate for our investigative analysis, because, when all other factors are held constant, models with more compact latent spaces generally produce more invariant representations [20].
-Thus, for the detailed analysis and evaluations, we pick ALBERT  [^2] We used open source Hugging Face albert-base-v2 and albert-xxlarge-v2 (https://huggingface.co/albert).
-[23] as our primary target since it has the tightest latent space dimensionality.
-Finally, we hypothesize that exploring the formation of such analogs early in the language model, potentially at the input embedding layer, could open the door to the possibility of knowledge-reasoning separation.
-Thus, we focus our investigation on this layer.
-More details about the models can be found in Appendix A.
-4 Experiments and Discussions
-Concept formation involves the creation of abstract internal representations through the clustering of inputs with shared features.
-While numerous intrinsic clustering metrics exist to assess cluster formation quality, their application to our use case is limited by two factors:
-  • The formed clusters are not situated within Euclidean space [24, 8, 7], rendering geometric properties such as cluster distances inadequate indicators of concept well-formedness and distinctiveness.
-  • Concepts inherently possess a degree of vagueness [19], thus metrics like compactness or separation do not reliably reflect the quality of formed concepts.
-Consequently, to assess the quality of the formed concepts, we shift our focus to evaluating concept alignment, specifically examining both LM-LM and Human-LM alignment.
-LM-LM alignment allows for the generalization of our observations to other language models, while Human-LM alignment serves as an indicator of the quality and real-world interpretability of the formed concepts.
-Concept Extraction Setup.
-We configured our algorithms to create $k$-NN graph iteratively for different values of $k$.
-This enables us to observe the communities/concepts formation at various granularities.
-Table 1 shows the number of identified clusters for $k=[100,75,50,25,12,6]$.
-Table 1: Number of communities with different granularities of nearest neighbors for ALBER, T5 and Glove. For Glove, we only used the subset of GloVe that present in Albert Vocab
-$K$ settings Vocab
-Models 100 75 50 25 12 6 Size
-T5 1 115 1203 4551 8137 9407 32000
-Albert 8 133 1058 4442 7718 8626 30000
-GloVe 9 207 1157 3521 6237 7200 25869
-4.1 LM-LM alignment
-To evaluate LM-LM alignment, we measure the community mapping across the language models under test.
-We define a mapping as the ratio of the successfully mapped communities from the embedding space of the language model with the smaller vocabulary to the language model with the larger vocabulary.
-Furthermore, we define success in mapping process as if the ratio of of found tokens from source community in the destination community is greater than a certain threshold.
-thresholds greater than 0.5 imply at worst many-to-one mapping (when the number of tokens in the source and destination space are similar, it is more likely to imply one-to-one mapping).
-Such one-to-one mapping implies a direct concept alignment, and many-to-one mapping implies the source community is aligned with hierarchies of the destination concept.
-This mapping score provides a quantitative measure of the degree to which the concept communities formed by one model can be translated to the other.
-To refine our analysis, we filtered the tokens to include only those that start with whitespace in T5 and ALBERT models.
-Empirically, this enhances the likelihood of associating the token with a word existing in the external world, thereby enhancing the validity of our concept alignment assessment.
-Note that to find the mappings we only used the token sets that are shared between the model pairs.
-That is 25869 for Albert/GloVe, 14861 for T5/GloVe and 5055 tokens for T5/Albert pairs.
-Table 2: Pair-wise Mapping score of the formed communities in the Glove, Albert, and T5 embedding spaces.
-Thre- Albert / GloVe T5 / GloVe T5 / Albert
-shold $K$=75 $K$=50 $K$=25 $K$=12 $K$=6 $K$=75 $K$=50 $K$=25 $K$=12 $K$=6 $K$=75 $K$=50 $K$=25 $K$=12 $K$=6
-0.5 0.70 0.57 0.62 0.68 0.70 0.62 0.74 0.84 0.89 0.91 0.61 0.78 0.91 0.95 0.97
-0.6 0.56 0.42 0.43 0.42 0.42 0.46 0.60 0.70 0.72 0.72 0.43 0.60 0.80 0.86 0.86
-0.7 0.42 0.30 0.30 0.28 0.27 0.27 0.46 0.54 0.56 0.56 0.30 0.49 0.70 0.79 0.80
-0.8 0.32 0.22 0.24 0.23 0.21 0.15 0.39 0.48 0.52 0.52 0.24 0.45 0.69 0.78 0.79
-0.9 0.17 0.16 0.20 0.21 0.21 0.07 0.32 0.45 0.50 0.51 0.18 0.43 0.67 0.78 0.79
-[…]
-The aforementioned observations strongly support the generalization of LM-LM concept alignment.
-The high degree of alignment, particularly for granular concepts and between models with similar architectures (like ALBERT and T5), indicates a shared understanding of semantic relationships across these models.
-This opens up exciting possibilities for leveraging this alignment to improve model performance or transfer knowledge between models without the need for extensive retraining.
-However, further research is needed to explore the extent and limitations of this potential knowledge transfer, especially for more abstract concepts and between models with different architectures.
-Case Sensitivity Analysis.
+Appendix F Models Under Investigation
+We used huggingface repository for all our models.
+F.1 Albert
+Albert (A Lite BERT) 23 is a transformer-based model for language representation learning, designed to be more efficient than its predecessor, BERT 11.
+While it shares the same basic architecture as BERT, it incorporates two main key modifications:
+  • Factorized embedding parameterization.
+The benefit of factorized embedding parameterization in Albert is the significant reduction in the number of parameters compared to models like BERT.
+In BERT, the word embedding size (E) is tied to the hidden layer size (H), leading to a large embedding matrix as H increases.
+Albert instead factorizes this embedding into two smaller matrices, one projecting token ID vectors to a lower-dimensional space (E) and another projecting from this space to the hidden layer (H).
+This allows H to be much larger than E without increasing the parameter count of the embedding layer substantially, resulting in a more efficient use of parameters.
+This is particularly beneficial for large models, where memory limitations can hinder training and deployment.
+  • Cross-layer parameter sharing.
+Parameter sharing acts as a form of regularization, preventing the model from overfitting to specific layers or features in the data.
+This can lead to improved generalization performance on unseen data.
+Note that reducing the number of parameters and sharing information across layers can force the model to learn more general representations, thus indirectly contributing to better conceptual groupings.
+datasets.
+Albert is pretrained on English text datasets, namely the English Wikipedia and BookCorpus, using self-supervised learning objectives.
+Tokenization. It uses Sentencepiece tokenizer 22 on the uncased corpus with a vocabulary size limit of 30K tokens.
+F.2 T5
+T5, or Text-to-Text Transfer Transformer, is a transformer-based architecture that casts all natural language processing (NLP) tasks into a text-to-text format.
+This means the model takes text as input and generates text as output, regardless of the specific task.
+At its core, T5 is an encoder-decoder model with the following key components:
+  • Encoder: This component takes the input text and processes it into a sequence of hidden representations.
+It uses multiple transformer layers, each consisting of self-attention mechanisms and feedforward neural networks.
+  • Decoder: This component generates the output text auto-regressively, conditioned on the encoder’s hidden representations.
+It also uses multiple transformer layers with self-attention and feedforward networks, as well as cross-attention mechanisms to attend to specific parts of the input sequence.
+Datasets. T5 is pre-trained on a massive dataset called C4 (Colossal Clean Crawled Corpus), which contains around 750 GB of clean English text.
+Tokenization. It uses Sentencepiece tokenizer on the cased corpus with a vocabulary size limit of 30K tokens.
+F.3 GloVe
+GloVe (Global Vectors for Word Representation) is a method for obtaining vector representations for words.
+Unlike context-based models like transformer-based LM models, GloVe leverages global word co-occurrence statistics across a corpus to learn word vectors.
+The GitHub repository [^5] https://github.com/stanfordnlp/GloVe provides an implementation of the GloVe model for learning word representations (word vectors or embeddings).
+We used the default embedding provided by the python API  [^6] Common Crawl (840B tokens, 2.2M vocab, cased, 300d vectors, 2.03 GB download), and a context window size of 10..
+Tokenization.
+They used Stanford tokenizer  [^7] https://nlp.stanford.edu/software/tokenizer.shtml, a form of BPE tokenization scheme 45, 13 that constructs unigram counts from a corpus, and optionally thresholds the resulting vocabulary based on total vocabulary size (2.2M 8882196016 cased tokens most frequent words in the case of GloVe embeddings) or minimum frequency count 40.
+F.4 Llama3
+LLaMA 3 (Large Language Model Meta AI) is a decoder-only large language model 1, which Grouped Query Attention (GQA) that allows the model to effectively handle longer contexts.
+The model is available in various sizes, including 1B, 3, and 70B parameters.
+The larger models exhibit significantly improved capabilities in reasoning and complex language tasks.
+Training Dataset.
+LLaMA 3 is pretrained on an extensive dataset, including over 15 trillion tokens sourced from diverse text corpora, such as books, articles, and websites.
+This large-scale training ensures comprehensive language understanding across different domains .
+Appendix G Case Sensitivity Analysis
 The T5 model’s vocabulary preserves case information, enabling us to examine how formed concepts align with capitalization differences.
@@
 We found that for highly granular concepts (k=6), 80% of these tokens belong to the same community.
-This ratio increases to 85% for k=25 before plateauing, suggesting that case variations generally do not significantly alter the semantic grouping of tokens.
+This ratio increases to 85% for k=25 before plateauing, suggesting that case variations generally do not drastically alter the semantic grouping of tokens.
 This finding supports the notion that the model learns to associate words with their meanings regardless of capitalization, particularly for more abstract or broader concepts (larger k values).
 However, the initial increase in alignment ratio with increasing k implies that case sensitivity might still play a minor role in differentiating highly specific or nuanced concepts.
-4.2 Human-LM alignment
-In the previous section, we established the concepts formed in LM models have a degree of alignment with respect to each other.
-To investigate human-LM alignment, in this section, we focus our analysis on the ALBERT model.
-While our empirical results are derived from ALBERT, the principles discussed are generalizable to other LLMs, given the LM-LM alignment established earlier (we include human-LM alignment evaluations for GloVe and T5 in Appendix B).
-Concept Hierarchy Formation.
-When examining the broader perspective (i.e., $k$=100), the model primarily found groups of named entities (mostly names of people and places), adverbs, sub-words, some number symbols, and a super-community for ‘other vocabulary’ (appendix C, Figure 4 shows the overall concept hierarchies of the ALBERT vocab).
-Zooming in further (e.g., $k$=75), these communities revealed more specific clusters that are relatable to the real world.
-For instance, within named entities, clusters formed for personal vs.
-location names, even further pinpointing locations by country.
-Similarly, within the aforementioned super-community, sub-clusters emerged that mapped to various lifestyle aspects, including relationships, social structures, multimedia, and motile creatures.
-As the granularity level increases (approaching a smaller $k$ value), clusters exhibit a stronger association with word forms.
-Concept Alignments. To investigate the presence and extent of human-LM concept alignment, we systematically examine the model’s representations of distinct facets of the external world:
-  • Numerical Concepts. This category focuses on numerical values, facilitating a targeted analysis of intra-category relationships, such as value-based ordering.
-  • Named Entities (Names and Locations).
-Primarily comprising proper nouns, this category leverages readily available labeled datasets to validate alignment between the LM’s conceptual groupings and established knowledge bases.
-It also allows us to investigate if associative relations (e.g., country-nationality) influence community formation.
-  • Social Structures.
-This category reflects the LM’s approach to forming categories that encompass both abstract and concrete nouns, including roles, titles, and processes.
-This analysis can elucidate its ability to grasp and align complex relationships and concepts.
-By analyzing these diverse facets, we aim to comprehensively evaluate the degree to which the LM’s internal representations align with human understanding of the external world.
-Table 3: Orderliness score of different communities of the numbers. Years refer to the cluster of numbers between 1816 - 2021. Support refers to the number of samples within the cluster.
-Category Top-1 Top-3 Top-5 Support
-0-100 0.68 0.89 0.92 100
-100-262 0.75 0.94 1.0 115
-263-300 0.46 0.78 0.89 37
-300-400 0.42 0.71 0.84 82
-400+ 0.53 0.75 0.83 110
-Years 0.86 0.96 0.98 203
-4.2.1 Symbols-Numbers
-At a high level, numbers and symbols are completely separated from the rest of the tokens in the embedding space.
-within themselves they form communities according to (i) years, (ii) integer values, (iii) tokens indicating monetary values (e.g., \$1), (iv) ratio/time (e.g., 2:1, 3:30), (v) fractions (e.g., ’/8’), (vi) large values (e.g., ‘100,000’), and (vii) percentages (e.g., $42\%$).
-Within these communities, integers form communities based on their hundreds.
-Within the embedding manifold, we analyzed the spatial distribution of number sequences, particularly focusing on their sequencing.
-Since the embedding manifold is not guaranteed to be a Euclidean structure [24, 8, 7], conventional distance-based measures of order were not applicable.
-We define a point $x$ as "locally ordered" on a manifold if it resides within the intersection of the $k$-nearest neighbor sets of both its immediate neighbors, $x+1$ and $x-1$.
-This localized approach captures the intuitive notion of orderliness within a small neighborhood on the manifold, rather than relying on a global ranking.
-Orderliness Score.
-Table 3 presents the orderliness scores of the communities identified by our algorithm.
-The "top-1" score represents strict orderliness, where $x+1$ lies exactly between $x$ and $x+2$ in the embedding space.
-The "top-3" and "top-5" scores measure relaxed orderliness, where $x+1$ falls within the 3 or 5 nearest neighbors of both $x$ and $x+2$, respectively.
-As the results indicate, all detected communities exhibit a higher degree of local orderliness, regardless of the chosen level of strictness.
-This finding is significant because it suggests that language models may possess not only the ability to categorize heterogeneous input entities but also the capacity to construct meaning within smaller, internally consistent structures (internally homogeneous sub-structures).
-This implies the potential for LMs to move beyond simply classifying information to actively interpreting and generating meaning within specific contexts.
-Figure 2: Visualization of the identified name and location communities of size larger than 10 entities. UMAP projection as well as kernel density estimate (KDE) plot is used for the visualization.
-4.2.2 Names and Locations
+Appendix H ALBERT Human-Model Aligment
+H.1 Names and Locations
 Figure 2 visualizes the major identified communities of locations and human names.
@@
 To mitigate the subjectivity risk of assessing the semantic structure, we further used external datasets in our evaluations.
-For Names, we used name-dataset [44] which consists of a comprehensive set of names (730K first names and 983K last names), their associated genders, and their popularity rank for each country.
-For locations, we used the country-state-city database [11] which contains information on 250 countries, 5K+ states, and 150K+ cities.
-Table 4(a) shows the high-level communities that our approach detected.
+For Names, we used name-dataset 43 which consists of a comprehensive set of names (730K first names and 983K last names), their associated genders, and their popularity rank for each country.
+For locations, we used the country-state-city database 12 which contains information on all countries, 5K+ states, and 150K+ cities.
+Table 9 shows the high-level communities that our approach detected.
 Most of the high-level communities are a mix of names/locations associated with certain geographical/cultural regions.
-Within these clusters, names and locations form distinct subclusters which we discuss in more detail in the following subsections.
-Table 4: Evaluation of the location and name tokens for different hierarchy levels.
-4(a) presents the precision for high-level communities, while 4(b) and 4(c) show the more granular communities formed within the detected communities in 4(a)
-(a) Precision of the largest identified name and location communities with respect to name and location databases.
+Within these clusters, names and locations form distinct sub-communities which we discuss in more detail in the following subsections.
+Note that since there’s no one-to-one mapping between names/locations in LLM vocabulary and external datasets (e.g., a name may appear in multiple countries, the external dataset is also a superset), recall is less relevant.
+We prioritize precision to evaluate accuracy, as it better reflects our ability to identify correct matches.
+Table 7: Precision of the largest identified name and location communities with respect to name and location databases.
 Note that these are at a higher level in the cluster hierarchies.
-Table 4(b) shows the identified granular sub-clusters and their associated precision.
+Table 9 shows the identified granular sub-clusters and their associated precision.
 Category Precision Support Note
-US/UK/AUS/NZ 0.882 1011 Human & Location Names
+US/UK/AUS/NZ 0.882 1011 Human & Location
 Male 0.854 946 Human Names
 Female 0.866 552 Human Names
-West-Asia 0.684 390 Human & Location Names
-Hispanic/Latino 0.685 282 Human & Location Names
+West-Asia 0.684 390 Human & Location
+Hispanic/Latino 0.685 282 Human & Location
 US 0.720 267 Location Names
-Europe 0.739 215 Human & Location Names
-East-Asia 0.741 178 Human & Location Names
-(b) The precision of communities based on the identified categories with respect to name-database. First indicate categories of first names last indicates last names
+Europe 0.739 215 Human & Location
+East-Asia 0.741 178 Human & Location
+Table 8: Precision of communities based on the identified categories with respect to name-database. First and last indicates category of first and last names.
 Country Overall Country Sup- Note
@@
 Japan 0.928 0.857 14 Last
-(c) The precision of communities based on the identified categories with respect to the location database.
+Table 9: Precision of communities based on the identified categories with respect to the location database.
 Country-Region Precision Support
 United States 0.80 240
-Germany 0.4125 80
+Germany 0.412 80
 France 0.409 66
-Africa 0.69 55
-India 0.58 50
+Africa 0.690 55
+India 0.580 50
 Italy 0.590 44
@@
 Spain 0.592 27
-China 0.5 20
+China 0.500 20
 Japan 0.736 19
-Philippines 0.46 15
+Philippines 0.460 15
 Pakistan 0.461 13
 Netherlands 0.636 11
-North-Africa 0.8 10
+North-Africa 0.800 10
 Names.
 To determine whether the sub-clusters are associated with names, we pre-filtered the clusters that at least 70% of their tokens are in the top 1000 names (of any country), with gender confidence of above 0.8.
-We primarily use gender confidence to distinguish between first-name from last name clusters.
+We primarily use gender confidence to distinguish between first-name from last-name clusters.
 Then we cross-referenced all identified name communities for all countries in the dataset and assigned the country name with the highest score  [^9] For some clusters we picked the second highest country name if the scores were similar.
@@
 as the cluster names.
-Table 4(b) shows the precision score of the identified granular communities.
+Table 9 shows the precision score of the identified granular communities.
 The first column shows the ratio of the community members that are indeed human names (overall precision), while the second column shows the ratio with respect to specific countries.
@@
 The reported precision score shows a high degree of categorization based on the country/culture of origin.
-It should be noted that we only included clusters with sizes larger than 10 and country-wise precision of larger than 0.5 due to space limitation.
+[^10] It should be noted that we only included clusters with sizes larger than 10 and country-wise precision of more than 0.5 due to space limitation.
 The list of identified cluster names goes far beyond the aforementioned table.
 Clusters such as character names from books, mythology, and car brand names were also identified which were not included due to space limitations.
+The formation of these clusters in the input embedding space, particularly those containing ethnic minority names, presents an opportunity for token engineering to mitigate potential ethnicity biases.
+(As our focus here is on interpretability and conceptual groupings and alignment, we provide an example of such a token engineering approach in Appendix 6 for interested readers.)
 Locations.
-It should be noted that, although the dataset is somewhat comprehensive, the LM token space associates only one token to each location, leading to an artificial decrease in precision for multi-token location names.
+It should be noted that, although the dataset is majorly comprehensive, the LLM token space associates only one token to each location, leading to an artificial decrease in precision for multi-token location names.
 Despite these limitations, we were able to identify communities within the input embedding space that are associated not only with the location category but also with specific regions/countries.
-The precision numbers in Table 4(c) suggest the model groups the locations with respect to the borders of the countries, which, to a certain degree, implies a subjective perception of geographical knowledge aligned with human perception (since it is evaluated against human annotated dataset).
-This finding suggests that the language model attempts to construct a human-like mental map  [^4] We refer the definition of the mental map to the [American Psychology Association dictionary] (https://dictionary.apa.org/mental-map).
+The precision numbers in Table 9 suggest that the model groups the locations with respect to the borders of the countries, which, to a certain degree, implies a subjective perception of geographical knowledge aligned with the external world  [^11] We refer the definition of the mental map to American Psychology Association dictionary.
 It is defined as “a mental representation of the world or some part of it based on subjective perceptions rather than objective geographical knowledge., wherein it approximates borders and associates nationalities with specific clusters/communities.
@@
 Note that dealing with multi-token location names is more challenging.
-For instance, "Carolina" was correctly clustered within the United States community by the LM, while our reference dataset misclassified it as a city in Brazil.
+For instance, "Carolina" was correctly clustered within the United States community by the LLM, while our reference dataset misclassified it as a city in Brazil.
 Additionally, not all location names with English spellings are included in the dataset; for example, ’Wurttemberg’ (or ’Nuremberg’), a region within Germany, is absent, leading to an artificial decrease in precision.
-4.2.3 Social Structures
+H.2 Social Structures
 Our methodology identifies a cluster of 903 members with the theme of social structures.
-In order to have a reference dataset, we annotated the dataset using a combination of GPT-4 and human annotators 666 A human annotator was used alongside GPT-4 to correct GPT-4 misclassifications.
+In order to have a reference dataset, we annotated the dataset using a combination of GPT-4 and human annotators 131313 A human annotator was used alongside GPT-4 to correct GPT-4 misclassifications.
 Label correction by a human happens faster than the label suggestion task.
 We estimated our approach is more time/cost-effective while resulting in the same quality annotations..
-Since analyzing such concepts can accompany a degree of human subjectivity, we feed the datasets to GPT-4 and ask to identify the theme of each category (as shown in Table 5) for mitigating such uncertainity.
+Given the potential human subjectivity in analyzing these concepts, we ask GPT-4 to identify the theme of each category in the datasets.
+(as shown in Table 10) for mitigating such uncertainity.
 In order to reduce subjectivity (we formed the class names based on the GPT-4 recommendation.
-Then, we asked GPT-4 and human annotator to classify each word with respect to given class names (we added another class named “Other" to avoid forcing the annotators to misclassification).
+Then, we asked GPT-4 and human annotator to classify each word with respect to given class names (we added another class named "Other" to avoid forcing the annotators to miscalassify).
 Then, we calculated the precision score of the identified clusters with respect to our annotated dataset.
@@
 Figure 4: Visualization of the social structure cluster and its associated identified sub-clusters.
-4.3 Knowledge-Reasoning Separation
-Thus far, we have demonstrated that the LM constructs a knowledge base (mental representation) directly within its input embedding layer.
-Furthermore, we have established a degree of human-LM alignment in both hierarchical structure and semantic meaning.
-Table 6: Huggingface ALBERT base model on GLUE, SuperGLUE, and SQUAD tasks.
-For the baseline, the model was finetuned without altering the embeddings.
-For the mid-point, the embedding layer entries are assigned the mid-point embedding of their associated community.
-We used [41] repo for benchmarking.
-mid-point GLUE
-Method/Tasks mnli mrpc qnli qqp rte sst stsb wnli
-baseline 0.827 0.841 0.902 0.858 0.765 0.915 0.872 0.549
-mid-point 0.849 0.865 0.910 0.875 0.783 0.922 0.890 0.563
-mid-point SuperGLUE SQUAD SQUAD
-Method/Tasks boolq cb copa multirc record wic wsc v1 (f1) v2 (f1)
-baseline 0.622 0.512 0.59 0.350 0.586 0.595 0.528 83.72 70.9
-mid-point 0.621 0.478 0.55 0.372 0.588 0.626 0.634 84.4 74.9
-Now, we are interested to see the extent to which the knowledge learned during pretraining is modular and separable from the reasoning mechanisms employed by the language model in downstream tasks.
-For example, can the knowledge learned during the pretraining phase be selectively removed or modified without significantly impacting the model’s performance on finetuning?
-The modularity can also impact the effectiveness of Language Model Inversion [35, 49] techniques, which aim to extract private information such as names or other sensitive details learned during the pre-training777 For example, if knowledge is found to be highly modular, it may be possible to develop targeted interventions that obscure or remove specific sensitive information without significantly impacting the model’s overall performance on downstream tasks..
-To investigate this, we selected GLUE, SuperGLUE, and SQUAD benchmarks as downstream tasks to assess language model performance.
-We then systematically removed within-community information by calculating and assigning the embedding space center (mid-point value) of each community to all its members.
-For example, in a community of names like "James," "John," and "Alex," all members would share the same embedding.
-Table 6 shows the results on major LM benchmarks GLUE, SuperGLUE, and SQUAD for the ALBERT base model.
-Although our experiment does not prove the separation of knowledge and reasoning, it indicates that at least the granular information acquired during the pretraining is not required for the model’s performance on the aforementioned LM benchmarks.
-This is significant because it opens the door for embedding engineering of private or harmful information that is learned during the pretraining.
-5 Limitations and Risks
-Limitations: The model forms the conceptual communities that are meaningful but its priority is not exactly the same as that of humans.
-The model vocabulary is a contributing factor to the way the model prioritizes the formation of conceptual clusters in its embedding layer.
-For example, the number of English names is much higher than the other languages and this has caused the model to form high-level communities (e.g., $k$=75) specified for names vs.
-less frequent names/locations a high-level community contains the combination of regions personal and location names.
-This limits our method to associate the KNN resolution with the abstraction level of the extracted concepts/categories.
-Risks: This work provides detailed information about (1) the formed clusters/concept in the input embedding layer, and (2) the separation of memory from reasoning in ALBERT.
-As the methodology can also be applied to other models, it can potentially facilitate more advanced adversarial attacks and content manipulation in language models.
-6 Conclusion and Future Works
-This research focused on concept formation and alignment across different language models (LMs) and between LMs and humans.
-We propose a concept extraction mechanism that uncovers the emergence of distinct conceptual communities within the entire embedding space.
-This approach gives a more holistic view of conceptual representation than examining specific samples, as it explores the full landscape of concepts embedded within the models.
-We examined GloVe, ALBERT, and T5 LMs and found that all of them form communities that can be aligned with the external world categories/concepts.
-We further investigated various aspects of the emergent conceptual communities, ranging from their numerical sequencing to the interrelationships and associations found within the blend of concrete and abstract concepts.
-Across various language models and types of concepts, distinct communities emerge in the embedding space.
-These communities demonstrate a significant alignment with human-understandable concepts, suggesting that statistical patterns, rather than specific architectures, or conceptual domains, drive concept formation across LMs.
-However, the quality and alignment of the formed concepts is a function of the architectural choices, training data, and tokenization schemas.
-We further observed that concept formation in transformer-based LMs occurs as early as the input embedding space.
-This suggests that the input embeddings themselves encode substantial semantic information, contrary to traditional views where meaning emerges later.
-Moreover, the emergent concepts within the input embedding layer exhibit a degree of separation from subsequent transformer layers.
-Early concept formation has several implications, opening exciting avenues for future line of works:
-  • Knowledge transfer across tasks and models: By identifying and leveraging shared conceptual structures across different tasks and models, early concept formation can enable more efficient transfer learning and reduce the need for extensive retraining on new tasks.
-  • Improving model architectures, training, and evaluation: Understanding how concepts emerge early in the model can inform the design of architectures that better capture and represent meaning, leading to improved performance and interpretability.
-This knowledge can also guide the development of more effective training techniques and evaluation metrics.
-  • Mitigating training data information leakages and reducing biases: Early concept formation can be utilized to identify potential biases and information leakages in the input embedding layer.
-This opens the door to the development of token embedding engineering techniques to mitigate these biases without requiring re-train.
-Appendix A Models Under Investigation
-A.1 ALBERT
-ALBERT (A Lite BERT) [23] is a transformer-based model for language representation learning, designed to be more efficient than its predecessor, BERT [10].
-While it shares the same basic architecture as BERT, it incorporates two main key modifications:
-  • Factorized embedding parameterization.
-The benefit of factorized embedding parameterization in ALBERT is the significant reduction in the number of parameters compared to models like BERT.
-In BERT, the word embedding size (E) is tied to the hidden layer size (H), leading to a large embedding matrix as H increases.
-ALBERT instead factorizes this embedding into two smaller matrices, one projecting token ID vectors to a lower-dimensional space (E) and another projecting from this space to the hidden layer (H).
-This allows H to be much larger than E without increasing the parameter count of the embedding layer substantially, resulting in a more efficient use of parameters.
-This is particularly beneficial for large models, where memory limitations can hinder training and deployment.
-  • Cross-layer parameter sharing.
-Parameter sharing acts as a form of regularization, preventing the model from overfitting to specific layers or features in the data.
-This can lead to improved generalization performance on unseen data.
-Note that reducing the number of parameters and sharing information across layers can force the model to learn more general representations, thus indirectly contributing to better concept formation.
-datasets.
-ALBERT is pretrained on English text datasets, namely the English Wikipedia and BookCorpus, using self-supervised learning objectives.
-Tokenization. It uses Sentencepiece tokenizer [22] on the uncased corpus with a vocabulary size limit of 30K tokens.
-A.2 T5
-T5, or Text-to-Text Transfer Transformer, is a transformer-based architecture that casts all natural language processing (NLP) tasks into a text-to-text format.
-This means the model takes text as input and generates text as output, regardless of the specific task.
-At its core, T5 is an encoder-decoder model with the following key components:
-  • Encoder: This component takes the input text and processes it into a sequence of hidden representations.
-It uses multiple transformer layers, each consisting of self-attention mechanisms and feedforward neural networks.
-  • Decoder: This component generates the output text auto-regressively, conditioned on the encoder’s hidden representations.
-It also uses multiple transformer layers with self-attention and feedforward networks, as well as cross-attention mechanisms to attend to specific parts of the input sequence.
-Datasets. T5 is pre-trained on a massive dataset called C4 (Colossal Clean Crawled Corpus), which contains around 750 GB of clean English text.
-Tokenization. It uses Sentencepiece tokenizer on the cased corpus with a vocabulary size limit of 30K tokens.
-A.3 GloVe
-GloVe (Global Vectors for Word Representation) is a method for obtaining vector representations for words.
-Unlike context-based models like transformer-based LM models, GloVe leverages global word co-occurrence statistics across a corpus to learn word vectors.
-The GitHub repository [^8] https://github.com/stanfordnlp/GloVe provides an implementation of the GloVe model for learning word representations (word vectors or embeddings).
-We used the default embedding provided by the python API  [^9] Common Crawl (840B tokens, 2.2M vocab, cased, 300d vectors, 2.03 GB download), and a context window size of 10..
-Tokenization.
-They used Stanford tokenizer  [^10] https://nlp.stanford.edu/software/tokenizer.shtml, a form of BPE tokenization scheme [47, 12] that constructs unigram counts from a corpus, and optionally thresholds the resulting vocabulary based on total vocabulary size (2.2M 1111112196016 cased tokens most frequent words in the case of GloVe embeddings) or minimum frequency count [40].
-Appendix B Human-LM Alignments For Glove, T5
-We performed our concept extraction algorithm on Glove, ALBERT, and T5 models.
-We observed that concept formation happens in all the models, however, the quality of the formed concepts is better in ALBERT model.
-Glove embeddings contain more than 200K tokens, in order to enable apple-to-apple comparison with ALBERT, we found the intersection of these tokens with the huggingface Albert-base-v2 token set, and, applied our concept extraction algorithm.
-We the formed clusters have weak correlation with the concepts formed in ALBERT model 121212if we assume Glove-ALBERT mapping exists between two clusters if more than half of their members are equivalent, %69 percent of the Glove clusters have a corresponding ALBERT cluster for K=6.
-(they similarities are stronger for concrete names/entities).
-Tables 7 and 8 show the T5 location and name communities detected by our algorithm.
-High precision numbers for these cluster indicate clear concept formation.
-However,since the pretrained HuggingFace T5 uses cased token set for the pretraining, the number of tokens in associated with location and names are much smaller than T5 and Glove.
-Tables 9 and 10 show the GloVe location and name communities detected by our algorithm.
-Although, it shows the categories are formed in GloVe embeddings as well, the numbers suggests the quality of the formed categories have less quality than both ALBERT and T5 counterpart.
-Table 7: Location communities founds in T5 tokens. Note that most of the communities are member of 0_0_5, and 0_0_13 super communities.
+Appendix I Human-LM Alignments For Glove, T5, Llama
+Table 11: Location communities founds in T5 tokens. Note that most of the communities are member of 0_0_5, and 0_0_13 super communities. SOAM stands for South America.
 Precision Support Cluster Name
@@
 Britain/Ireland 0.843 118 0_0_5_5
-Africa/South-America 0.672 64 0_0_13_2
+Africa/SOAM 0.672 64 0_0_13_2
 Germany 0.619 21 0_0_13_0_0
@@
 Baltic 1.000 4 0_0_13_1_8
+We performed our concept extraction algorithm on Glove, Albert, and T5 models.
+We observed that conceptual groupings happens in all the models, however, the quality of the formed concepts is better in Albert model.
+Glove embeddings contain more than 200K tokens, in order to enable apple-to-apple comparison with Albert, we found the intersection of these tokens with the huggingface Albert-base-v2 token set, and, applied our concept extraction algorithm.
+We the formed clusters have weak correlation with the concepts formed in Albert model 141414if we assume Glove-Albert mapping exists between two clusters if more than half of their members are equivalent, %69 percent of the Glove clusters have a corresponding Albert cluster for K=6.
+(they similarities are stronger for concrete names/entities).
+Tables 11 and 12 show the T5 location and name communities detected by our algorithm.
+High precision numbers for these cluster indicate clear conceptual groupings.
+However,since the pretrained HuggingFace T5 uses cased token set for the pretraining, the number of tokens in associated with location and names are much smaller than Albert and Glove.
+Table 13 shows the Llama3 name communities detected by our algorithm.
+Tables 14 and 15 show the GloVe location and name communities detected by our algorithm.
+Although, it shows the categories are formed in GloVe embeddings as well, the numbers suggests the quality of the formed categories have less quality than both Albert and T5 counterpart.
 Table 12: Name communities founds in T5 tokens. Note that most of the communities are member of 0_0_5 super community.
@@
 Male 0.846 Russian/Italy 0.615 13 0_0_5_15
-Politician Names 1.0 N/A N/A 14 0_0_5_13
+Politicians 1.0 N/A N/A 14 0_0_5_13
+Table 13: Examples of Name Entity clusters with size greater than 25 tokens found in Llama3 input embedding size. Note that we only evaluate against external datasets with English named entities.
+Type Precision Support Note Cluster names
+female 0.757 393 US/UK 0_0_0_0_0_0
+male 0.811 291 US/UK 0_0_0_0_0_2
+male 0.6639 171 US/UK 0_0_0_0_0_3
+male 0.495 105 Saudi/UAE 0_0_0_0_0_5
+male 0.6875 53 US/South Africa 0_0_0_0_0_11
+male 0.5588 46 UK/Canada 0_0_0_0_0_12
+female 0.571 41 US/UK 0_0_0_0_0_13
+male 0.551 37 Mexico 0_0_0_0_0_14
+male 0.619 30 Netherlands/US 0_0_0_0_0_16
+male 0.666 30 US/UK 0_0_0_0_0_17
+male 0.653 27 Biblical 0_0_0_0_0_18
+last names 0.724 340 US/UK 0_0_0_0_0_1
+last names 0.528 135 Canada/US 0_0_0_0_0_4
+last names 0.509 65 UK/US 0_0_0_0_0_8
+last names 0.717 39 Mexico/Chile 0_0_0_0_6_0
+locations 0.75 485 Mostly American 0_0_0_8_0
 Table 14: Location communities founds in the subset Glove tokens that exists in Albert Vocab.
@@
 UAE 0.6 0.467 no-gender 15 0_1_6_2
-Appendix C Extracted Concept Hierarchies
+Appendix J Extracted Concept Hierarchies
 Figure 5 shows the overall structure of hierarchical communities extracted by our proposed method.